### v1

# IF-Critic: Towards a Fine-Grained LLM Critic for Instruction-Following Evaluation

Abstract
Instruction following is a fundamental ability of Large Language Models (LLMs), requiring their generated outputs to follow multiple constraints imposed in input instructions. Numerous studies have attempted to enhance this ability through preference optimization or reinforcement learning based on reward signals from LLM-as-a-Judge. However, existing evaluation models for instruction following still possess many deficiencies, such as substantial costs and unreliable assessments. To this end, we propose IF-Critic, an LLM critic that can provide efficient and reliable assessments of constraint following in the instructions. We first develop a checklist generator to decompose instructions and generate constraint checklists. With the assistance of the checklists, we collect high-quality critique training data through a multi-stage critique filtering mechanism and employ a constraint-level preference optimization method to train IF-Critic. Extensive experiments demonstrate that the evaluation performance of IF-Critic can beat strong LLM-as-a-Judge baselines, including Deepseek-R1 and o4-mini. With the scalable reward signals provided by IF-Critic, LLMs can achieve substantial performance gains in instruction-following optimization under lower computational overhead compared to strong LLM critic baselines.
Work done when this author interned at Zhipu AI.Equal contribution.Corresponding author.
Codehttps://github.com/thu-coai/IF-CRITIC
Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in various NLP tasks (Zhao et al. 2023). Among these, instruction following stands out as one of the most crucial requirements for LLM applications (Ouyang et al. 2022). In the real-world use of LLMs, nearly all tasks are formulated as instruction following, where human instructions specify task requirements and impose corresponding constraints on the model output (Jiang et al. 2024). Accurately following instructions serves as a key factor in ensuring the helpfulness and reliability of LLMs (Huang et al. 2024).
However, LLMs still have not excelled in instruction following, especially when dealing with complex instructions that involve numerous constraints (Jiang et al. 2024; Wen et al. 2024). Numerous research attempts have been made to improve the instruction-following ability of LLMs. They often leverage LLMs to synthesize complex instructions and perform preference optimization or reinforcement learning based on the reward signals from LLM-as-a-Judge (Sun et al. 2024; Zhang et al. 2024c; He et al. 2024a; Ren et al. 2025; Liu et al. 2025b; Peng et al. 2025).
Figure 1: A usage example of IF-Critic: Given an instruction and a response, a checklist generator first decomposes the instruction to generate a constraint checklist. Then, IF-Critic can provide fine-grained evaluations for the response with respect to its following of all included constraints.
Although these works have achieved promising results, the usage of LLM-as-a-Judge for instruction-following still faces two major yet underexplored challenges: (1) Substantial Costs: Existing methods typically rely on proprietary LLMs (e.g., GPT-4o) or large reasoning models (e.g., QwQ-32B) to provide evaluation results as reward signals, which is costly and difficult to scale up in training data construction. (2) Unreliable Assessments: Current LLMs still struggle to accurately assess instruction-following (Zhang et al. 2024b), exhibiting low recall in error detection (Kamoi et al. 2024) and limited counting ability (Zhang, Cao, and You 2024; Fu et al. 2024). These limitations introduce noise to the evaluation and restrict the gains in model optimization. To tackle the above challenges, some works attempt to introduce handcrafted constraints that can be automatically verified by code execution in data construction, aiming to enhance the efficiency and reliability of evaluation (Dong et al. 2025; Kim et al. 2025; Liu et al. 2025b; Peng et al. 2025). However, these constraints are often homogeneous and atomized, failing to cover the diversity and complexity of human instructions, such as constraints composition (Wen et al. 2024).
In this work, we propose IF-Critic, an LLM critic with 14B parameters tailored for instruction-following evaluation. To enhance inference efficiency and obtain a more holistic and granular perception of the instructions, IF-Critic adopts a novel checklist-guided critique generation paradigm. As shown in Figure 1, we first utilize a checklist generator to decompose the instructions and generate constraint checklists. Then, IF-Critic can evaluate the following of all included constraints in one inference pass, eliminating the need to generate separate judgments for different constraints multiple times. To train IF-Critic, we elaborately prompt Deepseek-R1 (Guo et al. 2025a) to generate expert critiques guided by the checklists, then employ a multi-stage critique filtering mechanism to select the highest-quality critique for each constraint, which mitigates noise and significantly enhances the reliability of critiques for various constraints, including those involving counting. After fine-tuning on this data, we leverage the property that the critiques contain multi-constraint evaluation results in the checklist, further improving IF-Critic via a novel constraint-level preference optimization method. This method performs fine-grained, constraint-level comparisons between positive and negative critiques using Direct Preference Optimization (DPO) (Rafailov et al. 2023), reinforcing the perception of crucial preference information and leading to better alignment between IF-Critic and the expert critiques. Empirically, IF-Critic consistently outperforms various LLM-as-a-Judge baselines on four instruction-following meta-evaluation benchmarks, including the latest state-of-the-art LLMs Deepseek-R1 and o4-mini (Jaech et al. 2024).
Leveraging the fine-grained evaluation results from IF-Critic as reward signals, we adapt DPO and Group Relative Policy Optimization (GRPO) (Shao et al. 2024) to improve the instruction-following ability of LLMs. Experiments demonstrate that IF-Critic brings substantially greater improvements over strong baselines with significantly less computational overhead, highlighting its effectiveness and practicability. Our contributions can be summarized as follows:
  • We propose a novel framework for the development of instruction-following critics, including the checklist-guided critique generation paradigm, the multi-stage critique filtering mechanism for high-quality training data collection, and the constraint-level preference optimization method for model training.
  • We introduce IF-Critic, the first instruction-following critic with 14B parameters, developed through our framework. Experiments show that IF-Critic achieves superior evaluation performance compared to strong LLM-as-a-Judge baselines, including Deepseek-R1 and o4-mini.
  • We validate the potential of IF-Critic to provide scalable reward signals and improve the instruction-following abilities of LLMs using DPO and GRPO. IF-Critic can bring substantially greater improvements under lower computational overhead compared to strong baselines.
Related Work
LLM-based Evaluation.
Utilizing LLMs as text evaluators has gradually become prevalent (Wang et al. 2023a; Chen et al. 2023; Ke et al. 2023) due to their advantages in flexibility, interpretability, and generalization. Although proprietary LLMs have demonstrated excellent evaluation capabilities (Zheng et al. 2023; Wang et al. 2023a), their practical application is limited by issues such as high cost and potential data leakage (Ke et al. 2024). To reduce the reliance on proprietary LLMs, a series of works attempt to train smaller LLMs specialized in evaluation using the data generated from humans or proprietary LLMs (Kim et al. 2024a; Li et al. 2024; Ke et al. 2024; Hu et al. 2024; Kim et al. 2024b). Distinct from these works that focus on evaluating overall text quality, we develop an evaluation model that provides constraint-level, fine-grained assessments of instruction-following and leverages them as reward signals to enhance the instruction-following ability of LLMs.
Instruction Following Optimization.
Instruction following remains a critical factor in determining the practicality of LLMs (Liu et al. 2023; Lou, Zhang, and Yin 2024), driving extensive studies to enhance this ability. Earlier works primarily focus on following simple instructions that mostly consist of a single constraint (Chiang et al. 2023; Wang et al. 2023c). With the expanding application of LLMs, recent studies place greater emphasis on complex instruction-following that involves numerous constraints. They typically leverage LLMs to synthesize complex instructions and perform preference optimization or reinforcement learning based on the reward signals from LLM-as-a-Judge (Sun et al. 2024; Zhang et al. 2024c; Cheng et al. 2025; Ren et al. 2025; Liu et al. 2025b; Peng et al. 2025). Specifically, SPaR (Cheng et al. 2025) employs self-play tree search refinement through a fine-tuned refiner to construct DPO training data. RECAST (Liu et al. 2025b) categorizes constraint synthesis into soft and hard constraints, leveraging GPT-4o and code execution, respectively, to obtain evaluation results for GRPO training. However, the effectiveness of existing methods is restricted by the high costs and limited instruction-following evaluation ability of LLM-as-a-Judge, and code-verifiable constraints fail to cover the diversity and complexity of human instructions.
IF-Critic Framework
Figure 2: The pipeline of IF-Critic development. The left section illustrates the process of critique training data construction, while the right section presents the process of training IF-Critic.
Task Definition and Method Overview
Distinct from previous LLM critics that typically generate evaluation results for a single criterion or question (Wang et al. 2024a; Vu et al. 2024), IF-Critic leverages a constraint checklist to guide critique generation and evaluates the following of all constraints included in the checklist within one inference pass. Specifically, given an instruction $x$, corresponding response $y$, and a constraint checklist $\{c_{k}\}^{n}_{k=1}$ which contains all constraints within $x$, IF-Critic can provide a critique $C=\bigcup_{k=1}^{n}(e_{k},j_{k})$ for $y$, where each segment $(e_{k},j_{k})$ consists of an explanation $e_{k}$ and a binary judgment $j_{k}$ (0 or 1) for the respective constraint following. This paradigm not only enhances inference efficiency but also equips IF-Critic with a more holistic and granular perception of the instructions, thus improving the reliability of instruction-following assessments.
The development of IF-Critic consists of the following steps: After collecting complex user instructions and LLM-generated responses, we first develop a checklist generator to decompose the instructions and generate constraint checklists. With the assistance of the checklists, we elaborately prompt Deepseek-R1 to generate expert critiques and employ a multi-stage critique filtering mechanism to further enhance their quality. Finally, we utilize the collected expert critiques and perform a constraint-level performance optimization method to train IF-Critic. The overall pipeline is illustrated in Figure 2. IF-Critic can be leveraged to provide reliable reward signals, improving the instruction-following ability of LLMs via DPO or GRPO, as validated in our experiments.
Instruction and Response Collection
To obtain diverse instructions for critic training, we collect instructions from real application scenarios and utilize LLMs to categorize them according to the task taxonomy of CritiqueLLM (Ke et al. 2024), which encompasses 10 categories and covers diverse NLP applications in real-world scenarios. To improve data quality, we prompt LLMs to score instruction quality and develop a classifier to assess their constraint complexity, which is trained with a small amount of human annotation data. Finally, we balance the amount of instruction across task categories and acquire 55K high-quality instructions with relatively high complexity.
Then, we collect LLM responses from 15 representative models, which possess varying abilities in instruction-following, including GPT-4o (OpenAI 2023), GPT-4o-mini, seven versions of GLM-4 (GLM et al. 2024), three versions of Qwen2.5 (Yang et al. 2024), Deepseek-R1 (Guo et al. 2025a), Doubao1.5 (Doubao Team 2025), and Claude3.7-sonnet (Anthropic 2023). For each instruction, we randomly select two models to generate responses, resulting in a total of 110k evaluation inputs, each comprising an instruction and its corresponding LLM response.
Critique Training Data Construction
To assist IF-Critic in providing fine-grained assessments of instruction-following, we first develop a checklist generator to decompose the constraints within instruction $x$ and generate a constraint checklist $\{c_{k}\}^{n}_{k=1}$. We leverage a powerful LLM, Deepseek-R1111https://huggingface.co/deepseek-ai/DeepSeek-R1, for automatic checklist annotation in our collected instructions and utilize its generated checklists to fine-tune the checklist generator. Manual inspection of 1,000 test samples shows the accuracy of our checklist generator exceeds 99%, indicating its reliability.
With the assistance of the checklists, we proceed to employ Deepseek-R1 for expert critique annotation via a carefully designed prompt, which elicits a concise and specific explanation before the judgment for each constraint. Detailed prompts are in the Appendix. For each response, we collect $N$ critiques and adapt a multi-stage critique filtering mechanism to select the highest-quality critique for each constraint in the checklist, which is then used for training IF-Critic.
The mechanism comprises four stages. Specifically, we first employ cross-model and rule-augmented verification to filter out noise in critiques caused by two major limitations of LLM-as-a-Judge: potential evaluation bias and limited counting ability, respectively. Subsequently, we select the highest-quality judgements and explanations from the resulting noise-free critiques as the final critiques based on consistency, significantly enhancing their reliability.
Cross-Model Verification.
Previous works point out that LLM-as-a-Judge may suffer from potential biases (Zheng et al. 2023; Ye et al. 2025a), undermining the reliability of its generated critiques. To mitigate this, we utilize two additional LLMs, GLM-4-Plus and Qwen2.5-72B-Instruct, to independently verify two aspects of the expert critiques for each constraint: (1) The correctness of the explanation, and (2) The consistency between the explanation and judgment. Detailed prompts can be found in the Appendix. Any critiques that fail verification by any model in any aspect are strictly filtered, accounting for 11.3% of our data.
Rule-Augmented Verification.
Considering that LLMs often struggle with accurate counting (Zhang, Cao, and You 2024; Fu et al. 2024), we introduce rule-augmented verification to enhance the reliability of critiques involving length constraints (e.g., each title must be 8 characters long). Methodologically, we first prompt Qwen2.5-72B-Instruct to identify length constraints and extract the segments from the responses that are relevant to these constraints. Subsequently, we employ Python scripts to automatically identify length information within these segments, and require Deepseek-R1 to revise its critiques with the assistance of this information. The prompts for these steps are in the Appendix. This approach integrates the flexibility of LLM-based evaluation with the precision of rule-based evaluation, thereby effectively handling the diverse length constraints in human instructions.
Final Judgement Selection.
The correctness of judgment is vitally important in instruction-following evaluation. Given that critique generation is analogous to a Chain-of-Thought (CoT) (Wei et al. 2022) reasoning process, we incorporate self-consistency (Wang et al. 2023b), a widely used decoding strategy in CoT reasoning, to select final judgments and improve their correctness. Specifically, for each constraint $c_{k}$, we select the final judgment $j_{k}^{*}$ via majority voting among multiple expert critiques. Furthermore, previous works demonstrate that confidence, which is calculated as the proportion of the majority result to the total number of samplings, can serve as an important indicator of answer reliability (Xiong et al. 2024; Hu et al. 2024; Prasad et al. 2024), with higher confidence generally implying greater reliability. Inspired by this, we discard the final critiques of constraints with confidence lower than 0.75 in our experiment, which empirically balances data retention and reliability.
Final Explanation Selection.
After obtaining the correct final judgments, the last step is to select the highest-quality explanations that support these judgments. To achieve this, we devise a selection strategy inspired by Minimum Bayes Risk Decoding (Bickel and Doksum 2015). Specifically, for each constraint $c_{k}$, we first construct a hypothesis set $\mathcal{H}_{k}$ containing all explanations whose corresponding judgments align with the final judgments $j_{k}^{*}$. Then, we utilize a text similarity measure $u(\cdot)$ to calculate the consistency between each pair of explanations in $\mathcal{H}_{k}$. The explanation with the highest average consistency among others is selected as the final explanation $e_{k}^{*}$. The formal definition is as follows:
$\mathcal{H}_{k}=\{e_{k}^{(i)}\mid j_{k}^{(i)}=j_{k}^{*},i=1,2,…,N\}$ (1)
$e_{k}^{*}=\arg\max_{e\in\mathcal{H}_{k}}\frac{1}{|\mathcal{H}_{k}|}\sum_{\widetilde{e}\in\mathcal{H}_{k}}u(\widetilde{e},e)$ (2)
IF-Critic Training
After constructing the final critique training dataset $D=\{C_{i}=\bigcup_{k=1}^{n_{i}}(e_{ik}^{*},j_{ik}^{*})\}_{i=1}^{|D|}$, we partition $D$ into two subsets $D_{\text{sft}}\cup D_{\text{ref}}$ and employ a two-stage paradigm to train IF-Critic, comprising supervised fine-tuning (SFT) and constraint-level preference optimization. In the first stage, we utilize a critique generation prompt template $p$ to concatenate the instruction $x$, response $y$, and checklist $\{c_{k}\}^{n}_{k=1}$, applying a standard SFT loss on vanilla LLM $\theta$:
$\mathcal{L}_{\text{SFT}}=-\sum_{i=1}^{|D_{\text{sft}}|}\text{log}P_{\theta}(C_{i}|p_{i})$ (3)
After obtaining the SFT critic, we seek to further enhance its capability via preference optimization. Existing methods usually sample multiple independent responses from the policy model and utilize response-level rewards to select chosen and rejected examples (Yuan et al. 2024; Ye et al. 2025b; Yu et al. 2025). However, since the critiques in our task involve multiple segments for different constraints in the checklist, only the segments where chosen and rejected critiques differ in judgments encapsulate crucial preference information. Other segments may introduce unnecessary interference and dilute important supervision signals, leading to insufficient optimization (Chen et al. 2024; Cheng et al. 2025).
To mitigate this, we introduce fine-grained, constraint-level comparisons in preference pairs construction, thus reinforcing the perception of crucial preference information. Specifically, for each instruction $x$ in $D_{\text{ref}}$, we first sample $M$ critiques from the SFT critic. Then, we identify the best self-sampled explanation $\hat{e}_{k}$ for each constraint $c_{k}$ whose judgements align with the expert final judgement $j^{*}_{k}$ in a manner analogous to Equation 1 and 2 (if any). Finally, self-sampled critiques with at least one judgment that misaligns with the expert final critique are selected as rejected critiques $C_{l}=\bigcup_{k=1}^{n}(e_{k},j_{k})$. Retaining the segments that align with the expert final critique while replacing other misaligned segments with the best self-sampled explanations $\hat{e}_{k}$ to construct chosen critique $C_{w}$:
$(e^k,jk),jkjk\displaystyle C_{w}=\bigcup_{k=1}^{n}$ (4)
Rejected critiques are discarded when some misaligned segments lack substitute best self-sampled explanations. Consequently, we construct a preference dataset $D_{\text{dpo}}$ where the preference pairs only differ in the segments with inconsistent judgments, thereby reinforcing crucial preference information. Moreover, self-sampled critiques are closer to the SFT critic decoding space than the expert critique, which is generally beneficial for optimization (Guo et al. 2024). Finally, we conduct preference optimization via the following DPO loss:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=$ $-\mathbb{E}_{(p,C_{l},C_{w})\sim D_{\text{dpo}}}[\log\sigma(\beta\log\frac{\pi_{\theta}(C_{w}|p)}{\pi_{\text{ref}}(C_{w}|p)}$ (5)
$-\beta\log\frac{\pi_{\theta}(C_{l}|p)}{\pi_{\text{ref}}(C_{l}|p)})]$
where both $\pi_{\theta}$ (i.e., the trained policy model) and $\pi_{\text{ref}}$ (i.e., the fixed reference model) are initialized from the SFT critic.
Evaluation Results Utilization
Leveraging the evaluation results of IF-Critic as reward signals, we can apply DPO or GRPO algorithms to improve the instruction-following abilities of LLMs. For every input instruction $x$, the checklist generator produces the constraint checklist $\{c_{k}\}^{n}_{k=1}$, and the policy model $\pi_{\theta}$ samples $K$ candidate responses $\{y_{i}\}^{K}_{i=1}$. Then, the reward $r_{i}$ for each response $y_{i}$ can be determined by the generated critique $C_{i}=\bigcup_{k=1}^{n}(e_{ik},j_{ik})$ for $y_{i}$:
$r_{i}=\frac{1}{n}\sum_{k=1}^{n}j_{ik}$ (6)
Then, we can use this reward to construct DPO training data or to conduct GRPO training.
Experiments
Instruction-Following Evaluation
Evaluation Benchmarks.
Considering the scarcity of existing instruction-following meta-evaluation datasets, we sample instructions from four benchmarks: EvalCritic, CFBench (Zhang et al. 2024a), TRACE (Zhang et al. 2024c), and Multi-IF (He et al. 2024b) to construct meta-evaluation datasets. EvalCritic is a random split of the dataset we used for instruction-following optimization [^2] We elaborate on this dataset in the following ”Instruction-Following Optimization” section., whose distribution of instructions is distinct from the training data of IF-Critic. For the first three benchmarks, we randomly sample 150 instructions. For Multi-IF, we select the first round from the Chinese version, comprising 454 instructions. Responses for each instruction are generated by 2 out of 6 LLMs with varying capabilities. We use our checklist generator to generate constraint checklists for EvalCritic and directly adapt the original checklists for other benchmarks. For the first three benchmarks, two annotators are required to independently judge the following of each constraint, with a third inspector conducting cross-validation. Any discrepancies will be discussed to reach a consensus. For Multi-IF, we use the provided verification code for evaluation. Across these benchmarks, there are 1411 / 1070 / 1634 / 1220 constraints, respectively, of which 406 / 252 / 297 / 320 are labeled as ”not followed”.
Prompt Model EvalCritic CFBench TRACE Multi-IF (zh) Avg.
Positive Negative Positive Negative Positive Negative Positive Negative Average
F1 F1 F1 F1 F1 F1 F1 F1 F1
Constraint-Level o4-mini 0.887 0.717 0.927 0.755 0.923 0.599 0.960 0.890 0.832
Deepseek-R1 0.863 0.686 0.905 0.718 0.883 0.555 0.933 0.824 0.796
GPT-4.1 0.883 0.608 0.919 0.689 0.924 0.597 0.932 0.801 0.794
GLM-4-Plus 0.873 0.526 0.900 0.563 0.926 0.518 0.913 0.682 0.738
Qwen2.5-72B-Instruct 0.854 0.362 0.891 0.459 0.915 0.374 0.896 0.590 0.668
Llama-3.3-70B-Instruct 0.869 0.486 0.904 0.612 0.916 0.426 0.921 0.756 0.736
QwQ-32B 0.856 0.675 0.900 0.684 0.901 0.595 0.931 0.795 0.792
GLM-4-32B-0414 0.867 0.451 0.900 0.531 0.913 0.384 0.898 0.610 0.694
Checklist-Level o4-mini 0.911 0.753 0.935 0.761 0.925 0.639 0.964 0.900 0.849
Deepseek-R1 0.895 0.716 0.928 0.725 0.910 0.579 0.938 0.828 0.815
GPT-4.1 0.883 0.560 0.918 0.637 0.923 0.517 0.933 0.799 0.771
GLM-4-Plus 0.856 0.354 0.903 0.524 0.916 0.369 0.906 0.649 0.685
Qwen2.5-72B-Instruct 0.849 0.278 0.894 0.424 0.914 0.330 0.889 0.548 0.641
Llama-3.3-70B-Instruct 0.862 0.391 0.905 0.585 0.915 0.400 0.916 0.718 0.712
QwQ-32B 0.892 0.663 0.922 0.716 0.922 0.570 0.933 0.793 0.801
GLM-4-32B-0414 0.855 0.347 0.897 0.512 0.909 0.325 0.909 0.664 0.677
IF-Critic-14B 0.927 0.806 0.937 0.785 0.942 0.739 0.945 0.844 0.866
Table 1: Positive F1 and Negative F1 scores of constraint-following verification on the instruction-following evaluation task. The best performance among these models is bold, while the second-best performance is underlined.
Baselines.
We choose state-of-the-art general and evaluation-specific LLMs as our baselines. (1) General LLMs: We adopt 4 advanced proprietary LLMs, GLM-4-Plus (GLM et al. 2024), GPT-4.1 (OpenAI 2023), Deepseek-R1 (Guo et al. 2025a) and o4-mini (Jaech et al. 2024), as well as 4 strong open-source LLMs, GLM-4-32B-0414 (GLM et al. 2024), QwQ-32B (Team 2025), LLama-3.3-70B-Instruct (Grattafiori et al. 2024), and Qwen2.5-72B-Instruct (Yang et al. 2024) as our general baselines. These models can serve as evaluators for instruction-following via elaborate prompts without further training. We explore two prompting strategies: 1) Checklist-Level Prompt: requires the LLMs to evaluate the following of all constraints in the given checklist within a single inference pass, which is the same as IF-Critic. 2) Constraint-Level Prompt: requires the LLMs to evaluate each constraint individually. (2) Evaluation-specific LLMs: We select Skywork-Reward-V2-Llama-3.1-8B-40M (Liu et al. 2025a), Prometheus-BGB-8x7B (Kim et al. 2024b), RM-R1-Deepseek-Distilled-Qwen-32B (Chen et al. 2025), and RRM-32B (Guo et al. 2025b) as our evaluation-specific baselines. These models are mainly designed for pairwise comparison between two responses. To accommodate this evaluation setting, we convert the above datasets into pairwise versions. Specifically, for two responses generated from the same instruction, we select pairs in which one response follows all the constraints while the other does not to construct pairwise samples. The rewards determined by the critiques of IF-Critic according to Equation 6 for both responses are converted into pairwise comparisons to calculate the agreement rate. To ensure a fair comparison, we remove any samples that result in a tie for IF-Critic. Greedy search decoding is used for reproducibility.
Implementation Details.
We choose Qwen2.5-14B-Instruct (Yang et al. 2024) as the base model for both the checklist generator and IF-Critic. Deepseek-R1 is required to generate $N=5$ expert critiques per response. We use difflab [^3] https://docs.python.org/3/library/difflib.html to implement the text similarity measure $u(\cdot)$ for final explanation selection. During IF-Critic training, the ratio of $D_{\text{sft}}$ to $D_{\text{ref}}$ is set to 6:4. The number of self-sampled critiques, $M$, is set to 10, and up to 1 preference pair is constructed for each evaluation input. More details of IF-Critic training are in the Appendix.
Model EvalCritic CFBench TRACE Multi-IF
Skywork-Reward-V2-Llama-3.1-8B-40M 0.571 0.689 0.671 0.705
Prometheus-BGB-8x7B 0.464 0.556 0.529 0.489
RM-R1-Deepseek-Distilled-Qwen-32B 0.607 0.711 0.765 0.625
RRM-32B 0.534 0.689 0.765 0.636
IF-Critic-14B 0.964 0.956 0.882 0.977
Table 2: Pairwise agreement rates on the instruction-following evaluation task.
Method EvalCritic CFBench TRACE Multi-IF
IF-Critic-14B 0.861 0.863 0.840 0.895
Data Format
w/ Constraint-Level Critique 0.844 0.830 0.816 0.859
Critique Filtering
w/ Raw Data 0.814 0.792 0.774 0.780
w/o Cross-Model Verification 0.851 0.858 0.832 0.874
w/o Rule-Augmented Verification 0.827 0.823 0.789 0.825
w/o Final Judgement Selection 0.840 0.804 0.821 0.849
w/o Final Explanation Selection 0.840 0.846 0.807 0.858
Preference Learning
w/ Vanilla DPO 0.797 0.797 0.785 0.841
w/ Expert Critique 0.828 0.836 0.801 0.840
w/o Preference Learning 0.815 0.810 0.810 0.841
Table 3: The average of Positive F1 and Negative F1 under different ablation settings for IF-Critic.
Policy Model Method Reward Model / Critic Relative Multi-IF CFBench SysBench
Time Turn 1 Turn 2 Turn 3 CSR ISR PSR CSR ISR SSR
Qwen2.5-7B-Instruct - - - 76.14 61.65 51.41 0.79 0.46 0.56 60.33 35.72 19.10
Qwen2.5-32B-Instruct - - 83.02 72.43 63.12 0.84 0.55 0.65 81.76 66.44 44.83
Qwen2.5-7B-Instruct DPO Skywork-Reward-V2-Llama-3.1-8B-40M 0.79 77.86 63.12 53.15 0.82 0.53 0.63 61.59 38.60 23.60
QwQ-32B 13.4 80.44 66.05 56.82 0.82 0.50 0.61 64.85 41.32 24.23
IF-Critic-14B 1.00 81.25 67.51 57.73 0.83 0.52 0.63 67.40 44.84 28.71
GRPO Skywork-Reward-V2-Llama-3.1-8B-40M 1.04 75.28 61.12 51.90 0.81 0.50 0.60 65.82 44.48 25.79
QwQ-32B 3.08 78.59 63.51 52.60 0.83 0.54 0.64 76.25 59.28 37.58
IF-Critic-14B 1.00 81.87 67.36 58.76 0.85 0.59 0.69 81.19 66.80 44.44
LLama3.1-8B-Instruct - - - 73.13 64.27 56.84 0.70 0.32 0.43 60.14 37.32 19.38
LLama3.1-70B-Instruct - - 83.06 74.25 66.78 0.82 0.52 0.62 78.64 62.52 39.49
LLama3.1-8B-Instruct DPO Skywork-Reward-V2-Llama-3.1-8B-40M 0.79 76.71 69.04 60.59 0.75 0.38 0.50 67.88 47.53 28.22
QwQ-32B 12.8 79.72 71.94 62.97 0.76 0.41 0.52 68.36 46.97 27.53
IF-Critic-14B 1.00 80.75 72.52 64.13 0.76 0.41 0.53 69.67 49.04 28.15
GRPO Skywork-Reward-V2-Llama-3.1-8B-40M 1.33 75.57 67.88 58.59 0.77 0.42 0.53 76.99 59.96 37.42
QwQ-32B 3.10 80.67 72.49 64.40 0.77 0.44 0.53 78.22 62.32 40.79
IF-Critic-14B 1.00 83.83 73.32 65.05 0.81 0.50 0.59 79.90 66.40 44.27
Table 4: Experimental results (%) on Multi-IF, CFBench, and SysBench. Relative Time refers to the relative time to IF-Critic of reward calculation in DPO and average per-step training in GRPO. Skywork-Reward-V2-Llama-3.1-8B-40M is a reward model that generates scalar rewards, whereas QwQ-32B and IF-Critic generate natural language critiques.
Main Results.
Table 1 presents the F1-scores for both positive and negative classes in the instruction-following evaluation task of general LLMs, while Table 2 shows the pairwise agreement for evaluation-specific LLMs. Firstly, IF-Critic demonstrates superior performance on various evaluation benchmarks, outperforming all baselines on average F1. Notably, it surpasses the expert critiques generation models, Deepseek-R1, on all benchmarks, and o4-mini on three out of four benchmarks. In contrast, the performance of existing general LLMs remains limited, with relatively low negative F1. This aligns with our motivation for developing IF-Critic to enable more reliable evaluation. Secondly, existing evaluation-specific LLMs still perform poorly in evaluating instruction-following and lag far behind IF-Critic, highlighting the necessity of a critic specifically designed for this task. Finally, reasoning models (i.e., o4-mini, Deepseek-R1, and QwQ-32B) perform better with Checklist-Level Prompt, while other LLMs benefit more from Constraint-Level Prompt. We attribute this to two factors: (1) Evaluating multiple constraints at one inference pass increases task complexity. (2) With the assistance of the constraint checklist, long-chain reasoning enables a more holistic perception of the relationships between constraints. This validates the rationale of our checklist-guided critique generation paradigm.
Ablation Study.
We conduct ablation studies on three main factors of our framework: training data format, critique filtering mechanism, and preference optimization method. For the training data format, we decompose the expert critiques to train an alternative model that generates critiques for individual constraints. Table 3 shows that the performance of IF-Critic declines across all benchmarks, demonstrating that checklist-guided training can facilitate the perception of instructions and contribute to the final performance.
For the critique filtering mechanism, we remove each step individually to explore its impact on evaluation performance. Table 3 shows that all steps contribute to the final performance, with the removal of rule-augmented verification yielding the largest performance drop, underscoring its key role in improving the discriminability for length constraints. Notably, removing the entire mechanism and directly using all raw training data leads to the most severe performance decline, highlighting the effectiveness of critique filtering.
For the preference optimization method, we explored three variants of our method: (1) Constructing preference pairs directly based on the agreement rates between the expert critiques, denoted as Vanilla DPO; (2) Replacing misaligned segments of rejected critiques with expert critiques rather than the best self-sampled critiques; and (3) Performing SFT on the entire critique training dataset without preference optimization. Table 3 demonstrates that our constraint-level preference optimization method achieves the best performance, evidencing its effectiveness in enabling better alignment with expert critiques.
Figure 3: Explanation quality evaluation results. The percentages indicate the preference between IF-Critic and other evaluation models via human annotation.
Analysis on Explanation Quality.
To assess the quality of the explanations generated by IF-Critic, we randomly sample 300 constraints from the first three evaluation benchmarks and collect corresponding generated explanations from IF-Critic, QwQ-32B, Deepseek-R1, and o4-mini. For each pair of explanations (one from IF-Critic and the other from a baseline evaluation model, given the same evaluation input), we ask human annotators to judge which explanation is better (i.e., win, lose, or tie) in terms of correctness, informativeness, and helpfulness. The priority of these aspects follows the above order. As shown in Figure 3, IF-Critic surpasses QwQ-32B and Deepseek-R1 in win rates by 9.3% and 7.7%, respectively, and performs on par with o4-mini. This demonstrates that IF-Critic can generate high-quality explanations for instruction-following evaluation.
Instruction-Following Optimization
Training Datasets & Settings.
To collect diverse and high-quality complex instructions for model training, we first curate a set of seed instructions from real application scenarios. These scenarios are distinct from the training data of IF-Critic and generally exhibit lower instruction complexity. Stratified sampling based on task categories and LLM-based quality filtering is adopted to enhance their diversity and quality. Subsequently, we employ Deepseek-R1 to introduce constraints to these seed instructions, guided by a comprehensive constraint taxonomy encompassing 6 primary constraint categories and 4 constraint composition types, thereby increasing their complexity. Deepseek-R1 is also leveraged to validate the generated instructions, ensuring they are reasonable, unambiguous, and logically consistent. Further details are in the Appendix. In total, we obtain 18.5k high-quality complex instructions for training.
We choose Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct as the initial policy models for DPO and GRPO training. Three evaluation models are used to provide reward signals: IF-Critic, the state-of-the-art reward model Skywork-Reward-V2-Llama-3.1-8B-40M, and open-source general LLM QwQ-32B. For DPO training, we sample 10 responses per instruction, with a temperature of 1.0 and a top-p value of 0.9. The responses with the highest and lowest rewards are used to construct preference pairs. For GRPO training, we perform 32 rollouts per instruction for value estimation. More training details can be found in the Appendix.
Evaluation Benchmarks.
We evaluate our models on three popular instruction-following benchmarks: Multi-IF (He et al. 2024b), which assesses multi-turn and multilingual hard constraints following; CFBench (Zhang et al. 2024a), which covers a comprehensive range of constraint types; and Sysbench (Qin et al. 2025), which focuses on the following of system prompts. GPT-4o-2024-11-20 is used as the evaluation model for CFBench and SysBench. For Multi-IF, we follow RPO (Huang et al. 2025) and report the average performance across the English and Chinese subsets.
Main Results.
The main results are presented in Table 4. Firstly, optimization with IF-Critic brings substantially superior performance gains across various benchmarks compared to all strong baselines, substantiating its advanced capacity of reliable instruction-following evaluation. Notably, LLMs after GRPO training with IF-Critic even achieve comparable performance to significantly larger LLMs (32B or 70B) of the same model family. In contrast, the general reward model, Skywork-Reward-V2-Llama-3.1-8B-40M, is not well-suited for instruction-following optimization, yielding only marginal improvements. Secondly, compared to the powerful baseline LLM critic, QwQ-32B, IF-Critic only incurs less than a third of computational overhead, highlighting its practicality. Finally, between the two commonly used optimization methods, DPO and GRPO, GRPO proves more effective for instruction-following tasks, delivering greater performance gains. The reward curves during GRPO training are provided in the Appendix. We also validate that optimization with IF-Critic does not compromise the general performance of LLMs in the Appendix.
Conclusion
In this work, we propose IF-Critic, a LLM critic which is capable of providing fine-grained assessments for instruction-following. IF-Critic adopts a checklist-guided critique generation paradigm that evaluates the following of all constraints in a given checklist within one inference pass, improving its efficiency and reliability. To construct IF-Critic, we collect high-quality critique training data through a multiple-stage critique filtering mechanism and employ a constraint-level preference optimization method to train IF-Critic. Extensive experiments demonstrate that IF-Critic can beat strong LLM-as-a-Judge baselines and bring substantially superior performance gains to LLMs in instruction-following optimization, compared to strong LLM critic baselines.
Appendix A Details of IF-Critic Training
We use the Llama-Factory (Zheng et al. 2024) framework and the Zero Redundancy Optimizer (ZeRO) (Rajbhandari et al. 2020) stage 3 from the Deepspeed library (Rasley et al. 2020) to train IF-Critic. All experiments are conducted on 8 H800 GPUs. We utilize the AdamW (Kingma and Ba 2014) optimizer with a weight decay of 0.1, and set the maximum sequence length to 8192 tokens during all the experiments.
The overall API cost for critique training data collection is approximately 4000\$. During the supervised fine-tuning, the peak learning rate is set to 5e-6 with a 10% warmup ratio and a linear scheduler. The batch size is 16, and training is conducted for 2 epochs. During the constraint-level preference-optimization of IF-Critic, the peak learning rate is set to 1e-6 with a 10% warmup ratio and a cosine scheduler. The batch size is 128. Training utilizes a sigmoid loss function with a beta value of 0.1 and spans 1 epoch. Following previous works (Hou et al. 2024), an additional SFT loss is added to the chosen critique with a weight of 0.1.
Appendix B Details of Training Instructions Collection in Instruction-Following Optimization
To ensure the diversity and complexity of constraints, we first construct a comprehensive and hierarchical constraint taxonomy, which consists of 6 primary categories (i.e., format constraints, content constraints, behavioral constraints, style constraints, role constraints, and environmental constraints) and 54 secondary categories. During constraint synthesis, we randomly sample 2 to 5 secondary constraint categories and require Deepseek-R1 to introduce several corresponding constraints into the seed instructions, thereby increasing their complexity. In contrast to previous work that only considers independent and atomic constraints (Jiang et al. 2023; Ren et al. 2025; Peng et al. 2025), we further ask Deepseek-R1 to compose multiple constraints according to the composition types proposed in ComlexBench (Wen et al. 2024), including: (1) And: Coordination between different constraints, (2) Chain: Sequential completion of constraints, (3) Selection: Conditional selection of constraints, and (4) The nested structure of the above composition types. In this way, our constraint synthesis can more comprehensively encompass the complex instruction types in real-world scenarios.
In this scenario, it is notable that most of the hard constraints cannot be reliably verified using automatic verification codes alone, such as constraints targeting for specific segments of the response (e.g., Generate 10 titles, each no more than 8 words) and the composition of soft and hard constraints (e.g., Bold all adjectives that express feelings, such as **happy**), as it is challenging for verification code to accurately extract the corresponding segments of the response that should follow these constraints. We provide some examples of the seed instructions and the corresponding final constructed complex instructions in Table 5.
Type Prompt
Seed Instruction You’re an expert in artificial intelligence training. How is a machine learning model built? Please explain in a clear and easy-to-understand tone.
Final Instruction You’re an expert in artificial intelligence training. Please explain the process of building a machine learning model in five stages, following a “general-to-specific-to-general” structure. Ensure the following requirements are satisfied:
1. Use the symbols ❶❷❸ to notate each sub-step within each stage.
2. Do not include any mathematical formulas or code examples.
3. Integrate real-life kitchen or cooking analogies within each stage.
4. Finally, use $\triangleright$ to highlight three common mistakes people make when building a machine learning model.
5. Write in the second-person perspective, ensure no technical terms exceed middle school math knowledge, and make sure all analogies involve kitchen or cooking elements.
Seed Instruction Translate the following English words into Chinese:
attachment, surpasses, dissent, retention, unobtrusive, glazed, entails, confederates, skeptical, perceive, enormity, fanatical
Final Instruction Please translate the following English words into Chinese, strictly following these requirements:
1. Precede each translated Chinese word with a numbered bullet point (e.g., ①, ②, ③).
2. Indicate the part of speech of the original English word (e.g., noun/verb/adjective) in parentheses after each translation.
3. Provide a Chinese synonym after each Chinese translation word, separated by a dash.
Words to be translated:
attachment, surpasses, dissent, retention, unobtrusive, glazed, entails, confederates, skeptical, perceive, enormity, fanatical
Seed Instruction Write a conversation between a Chinese international student and a Japanese student who decide to try a new restaurant together. While they’re having their meal, one of them suggests that they go to see a movie afterward.
Final Instruction Please write a conversation between Xiaolin, a Chinese international student, and Yamada, a Japanese student. In the conversation, they decide to try a new restaurant together. While they’re having their meal, one of them suggests that they go to see a movie afterward. Ensure the following requirements are satisfied:
1. The conversation should contain three stages: (1) Stage 1: Arranging to dine at the restaurant (at least 2 turns), (2) Stage 2: Discussing the food during the meal (at least 3 turns), (3) Stage 3: Proposing to watch a movie (at least 3 turns). Separate each stage of the conversation with two blank lines.
2. If the restaurant is a Chinese restaurant, the suggested movie genre should be kung fu films. If the restaurant is a sushi restaurant, the suggested movie genre should be animated films.
3. Xiaolin should naturally mention Mid-Autumn Festival customs in the conversation, while Yamada should incorporate elements of the Obon Festival. Please use specific festival details (e.g., types of mooncakes or forms of Bon Odori).
4. Please write the conversation in Chinese. Indicate the speaker (Xiaolin: / Yamada:) at the start of each turn. The entire conversation should be no more than 10 turns.
Please first determine the restaurant type (Chinese restaurant or sushi restaurant), and then write the conversation according to the corresponding scene.
Table 5: Some examples of the seed instructions and the corresponding final constructed complex instruction in the instruction-following optimization.
Appendix C Details of Instruction-Following Optimization
For DPO training, we use the Llama-Factory (Zheng et al. 2024) framework, and all other settings are the same as the preference optimization for IF-Critic. During the reward calculation, IF-Critic and QwQ-32B utilize the vllm (Kwon et al. 2023) inference framework, while Skywork-Reward-V2-Llama-3.1-8B-40M is deployed by sglang [^4] https://github.com/sgl-project/sglang (since this reward model is not compatible with the vllm framework currently). The reward calculation and DPO training are both conducted on 8 H800 GPUs.
For GRPO training, we use the VeRL (Sheng et al. 2025) framework. The learning rate is set to 2e-6, with the maximum sequence length for both prompts and responses set to 4096 tokens. The train-batch size and PPO mini-batch size are each set to 32. For rollout responses generation, we employ the vllm (Kwon et al. 2023) framework, utilizing 60% of the available GPU memory, with a temperature and top-p value of 1.0. To encourage stable training, we also incorporate KL divergence regularization with a coefficient of 0.001, using the low-variance KL implementation. We save checkpoints every 30 steps during training. Following TULU 3 (Lambert et al. 2024) and VerIF (Peng et al. 2025), we use IFEval (Zhou et al. 2023) as the validation set to select the best checkpoint. All experiments are conducted on 24 H800 GPUs, with 8 GPUs dedicated to model training and the remaining 16 GPUs allocated for API deployment of the instruction-following critic or reward model. The deployment framework is identical to the inference framework described in the above DPO training.
Figure 4: Reward curves during GRPO training when IF-Critic and QwQ-32B are employed as the LLM critics. For LLama-3.1-8B-Instruct, training with QwQ-32B results in a model collapse after 300 steps, with the model tending to generate extensive repetitive and meaningless content. Due to efficiency considerations, we terminate further training and calculate the average per-step training time for all critics and reward models based on the first 300 steps.
Figure 5: Reward curves during GRPO training when Skywork-Reward-V2-Llama-3.1-8B-40M is employed as the reward model.
Policy Model Method Reward Model / Critic AlignBench MMLU-Pro Omni-MATH HumanEval Avgerage
Qwen2.5-7B-Instruct - - 58.61 57.32 32.20 79.27 56.85
DPO Skywork-Reward-V2 61.36 57.41 31.95 75.00 56.43(-0.42)
QwQ-32B 60.81 57.14 32.05 77.44 56.86(+0.01)
IF-Critic-14B 61.16 57.20 32.10 77.44 56.98(+0.13)
GRPO Skywork-Reward-V2 61.49 57.43 31.18 74.39 56.12(-0.73)
QwQ-32B 62.64 57.51 32.69 73.78 56.66(-0.19)
IF-Critic-14B 64.83 56.57 32.11 76.83 57.59(+0.74)
LLama3.1-8B-Instruct - - 40.66 44.17 16.05 64.02 41.23
DPO Skywork-Reward-V2 44.93 44.75 15.44 63.41 42.13(+0.90)
QwQ-32B 41.23 44.61 16.54 65.85 42.06(+0.83)
IF-Critic-14B 45.76 44.11 16.02 63.41 42.33(+1.10)
GRPO Skywork-Reward-V2 46.97 45.65 17.38 64.02 43.51(+2.28)
QwQ-32B 45.95 44.76 17.25 60.37 42.08(+0.85)
IF-Critic-14B 46.26 45.52 17.75 60.98 41.98(+1.40)
Table 6: Experimental results (%) on various general benchmarks. The original results of AlignBench range from 1 to 10; we convert the scale to 10-100 for ease of calculating average performance. Skywork-Reward-V2 refers to Skywork-Reward-V2-Llama-3.1-8B-40M.
Appendix D Reward Curves During GRPO Training
We provide the reward curves during GRPO training when IF-Critic and QwQ-32B are employed as the LLM critic in Figure 4, and Skywork-Reward-V2-Llama-3.1-8B-40M as the reward model in Figure 5. IF-Critic has a stronger error-detection capability in instruction following, which leads to it generally assigning lower rewards to policy models than QwQ-32B. We also observe that the rewards provided by Skywork-Reward-V2-Llama-3.1-8B-40M may have potential length bias. Using it as a reward model for GRPO training significantly increases the average response length of the policy model, resulting in a longer average per-step training time compared to IF-Critic.
Appendix E Experimental Results on General Benchmarks
We provide the performance of LLMs after instruction-following optimization in four general benchmarks: AlignBench (Liu et al. 2024), MMLU-Pro (Wang et al. 2024b), Omni-MATH (Gao et al. 2025), and HumanEval (Chen et al. 2021) in Table 6. GPT-4o-2024-11-20 is used as the evaluation model for AlignBench, and we use OmniJudge [^5] https://huggingface.co/KbsdJames/Omni-Judge for the evaluation of Omni-MATH. The results reveal that conducting instruction-following optimization with IF-Critic does not harm the general performance of LLMs.
Appendix F List of Prompt Templates
This section lists all the prompt templates applied throughout this work, including the prompt for scoring the quality of user instruction in Table 7, the prompt for constraint checklist generation in Table 8, the prompt for checklist-guided critique generation in Table 9, the prompt for cross-model verification in Table 10, and the prompt for rule-augmented verification in Table 11 and 12.
Appendix G Human Annotation Guideline for Constraint Verification
The human annotation guideline for constraint verification in the section Instruction-Following Evaluation is shown in Table 13. Given an instruction and a corresponding model response, as well as the constraint checklist for the instruction, human annotators are instructed to judge whether the model response follows each constraint in the checklist.
You are an expert specializing in evaluating the quality of user instructions. Please evaluate the quality of the user instruction based on the following criteria.
## Evaluation Criteria:
1. Low Quality: The user instruction contains significant issues, such as unclear, incomplete, or ambiguous information, making it impossible to determine the intent behind the instruction.
2. Medium Quality: The user instruction may include some unclear, incomplete, or ambiguous elements; however, the overall intent can still be inferred.
3. High Quality: The user instruction is clear, complete, and unambiguous, allowing the intent to be easily and definitively understood.
## Note:
1. You do not need to answer the user instruction, but only need to output the evaluation result.
2. If the user instruction requires additional retrieval or the use of tools to obtain an answer, it should be evaluated as Medium Quality.
Here are some examples and the user instruction to be evaluated:
[The Start of Examples]
{in_context_examples}
[The End of Examples]
[The Start of User Instruction]
{instruction}
[The End of User Instruction]
## Output Format
Prompt Quality: Low Quality / Medium Quality / High Quality
## Please directly output the evaluation result:
Table 7: The prompt template for scoring the quality of user instructions.
You are an information extraction expert specializing in identifying all constraints within instructions.
We will provide you with a user instruction. Your task is to extract all constraints present in that instruction. Constraints refer to all instructional content, excluding auxiliary information such as background knowledge, text materials, and in-context examples. This includes, but is not limited to, descriptions of basic tasks to be completed, and specific requirements on output content, style, and format. Output all constraints using this format exactly:
[The Start of Constraint 1]
Constraint: … (Please generate a specific constraint of the instruction. It must be complete and detailed, with no information omitted or altered in any way)
[The End of Constraint 1]
[The Start of Constraint 2]
Constraint: … (Please generate a specific constraint of the instruction. It must be complete and detailed, with no information omitted or altered in any way)
[The End of Constraint 2]
Remember the following points:
(1) Each constraint you output must be a verbatim excerpt from the given instruction. You must not fabricate, paraphrase, or add any content.
(2) You must extract and output all constraints in the order in which they appear in the instruction, without omitting any or inventing constraints not present in the instruction. To reiterate: Constraints refer to all instructional content, excluding auxiliary information such as background knowledge, text materials, and in-context examples.
(3) Do not output duplicate constraints. Each portion of the given instruction should appear in at most one constraint, and must not be repeated across multiple constraints.
(4) Each constraint should be atomic and independent, without inclusion or dependency relationships. At the same time, ensure appropriate granularity of constraints: neither too fine nor too coarse. Each constraint should have complete semantics and be understandable on its own, without reference to other constraints. At the same time, relatively independent constraints must not be merged into a single constraint.
Here are the user instruction:
[The Start of User Instruction]
{instruction}
[The End of User Instruction]
Table 8: The prompt template for constraint checklist generation.
You are a fair judge, specializing in evaluating the quality of an AI assistant’s responses to user instructions.
We will provide you with a user instruction, a constraint checklist for the user instruction, and an AI assistant’s response. Please think carefully and provide a detailed analysis of whether the AI assistant’s response follows each item on the constraint checklist. You need to analyze and assess every constraint in the checklist, and you do not need to analyze anything that is not listed in the checklist. You must output your explanation and judgment for each constraint, strictly following the format below:
[The Start of Constraint 1]
Constraint: … (Directly output the first constraint from the checklist here with no modifications)
Explanation: … (Provide a thorough, step-by-step analysis of whether the AI assistant’s response follows this constraint, referencing specific details of the response)
Judgment: [[The AI assistant’s response follows this constraint]] or [[The AI assistant’s response does not follow this constraint]]
[The End of Constraint 1]
[The Start of Constraint 2]
Constraint: … (Directly output the second constraint from the checklist here with no modifications)
Explanation: … (Provide a thorough, step-by-step analysis of whether the AI assistant’s response follows this constraint, referencing specific details of the response)
Judgment: [[The AI assistant’s response follows this constraint]] or [[The AI assistant’s response does not follow this constraint]]
[The End of Constraint 2]
Remember the following points:
(1) You should only focus on whether the response follows the items in the checklist, not on the overall quality of the response. Other shortcomings in the response (such as not fulfilling a constraint to a high standard) should not affect your judgment unless the constraint is not followed.
(2) Your judgment should be as strict as possible. You should output ”[[The AI assistant’s response follows this constraint]]” only if the response completely follows every part of the constraint as stated, without any omission or mistake.
(3) When analyzing a given constraint, you should not take into account whether other constraints are followed.
Here are the user instruction, constraint checklist, and AI assistant’s response:
[The Start of User Instruction]
{instruction}
[The End of User Instruction]
[The Start of Constraint Checklist]
{checklist}
[The End of Constraint Checklist]
[The Start of AI Assistant’s Response]
{response}
[The End of AI Assistant’s Response]
Table 9: The prompt template for checklist-guided critique generation.
Template Prompt
Correctness
Verification
You are a fair judge, specializing in evaluating the quality of an AI assistant’s responses to user instructions.
We will provide you with a user instruction, an AI assistant’s response, a specific constraint from the user instruction, and a critique of the AI’s response regarding whether the response follows this constraint from the user instruction. The critique comprises two parts: ”Explanation” and ”Judgment”. Your task is to carefully analyze this critique to judge whether it is reasonable and correct. After your analysis, output either ”[[The given critique is correct]]” or ”[[The given critique is not correct]]” as your final conclusion.
A correct critique must simultaneously satisfy all of the following three principles:
(1) Only focus on whether the response follows the constraint, not on the overall quality of the response. Other shortcomings in the response (such as not fulfilling the constraint to a high standard) should not affect the judgment unless the constraint is not followed.
(2) Be as strict as possible. Output ”[[The AI assistant’s response follows this requirement]]” only if the response completely follows every part of the requirement as stated, without any omission or mistake.
(3) Independently analyzes the given constraint, and does not take into account whether other constraints are followed.
Here are the user instruction, the constraint, the AI assistant’s response, and the critique to be judged:
[The Start of User Instruction]
{instruction}
[The End of User Instruction]
[The Start of Constraint]
{constraint}
[The End of Constraint]
[The Start of AI Assistant’s Response]
{response}
[The End of AI Assistant’s Response]
[The Start of Critique]
{critique}
[The End of Critique]
Consistency
Verification
You are a fair judge, specializing in evaluating the quality of an AI assistant’s responses to user instructions.
We will provide you with a constraint, a critique that comprises an explanation and a judgment generated by an evaluator, whose main purpose is to analyze whether an AI assistant’s response follows the constraint. You do not need to consider the content of the constraint or whether the critique is correct. You only need to judge whether the explanation and judgment within the critique are logically consistent, specifically:
(1) If the explanation section determines that the response follows the constraint, the judgment should be: ”[[The AI assistant’s response follows the constraint]]”.
(2) If the explanation section identifies any mistake or shortcoming in the response, indicating that the constraint is not fully followed, the judgment should be: ”[[The AI assistant’s response does not follow the constraint]]”.
Remember the following points:
(1) There should not be situations such as ”partially follow the constraint”. As long as the explanation identifies any mistakes or shortcomings in the response that result in the constraint not being fully followed, the judgment should be: ”[[The AI assistant’s response does not follow the constraint]].”
(2) If the explanation states that the constraints and the response are unrelated and there is no need for evaluation, the judgment should be: ”[[The AI assistant’s response follows the constraint]].”
Please analyze the logical consistency between the explanation and judgment following the above principles. Output your analysis first, and then output either ”[[Explanation and Judgment are consistent]]” or ”[[Explanation and Judgment are not consistent]]” to represent your final verdict.
Here are the constraint and the critique to be judged:
[The Start of Constraint]
{constraint}
[The End of Constraint]
[The Start of Critique]
{critique}
[The End of Critique]
Table 10: The prompt templates for cross-model verification.
You are a fair judge, specializing in evaluating the quality of an AI assistant’s responses to user instructions.
We will provide you with a user instruction, an AI assistant’s response, and a specific constraint from the user instruction. Please first determine whether the given constraint is related to length (e.g., ”The article should be no less than 800 words” or ”Each title must be 8 characters long”). If so, please extract the segments from the AI assistant’s response that should follow this length requirement (extract the original text from the response without any modifications).
## Reference Example
{in_context_examples}
Here are the user instruction, the constraint, and the AI assistant’s response:
[The Start of User Instruction]
{instruction}
[The End of User Instruction]
[The Start of Constraint]
{constraint}
[The End of Constraint]
[The Start of AI Assistant’s Response]
{response}
[The End of AI Assistant’s Response]
Remember the following points:
(1) The given constraints may contain length requirements for multiple segments of the response. Therefore, numerous segments of content may need to be extracted.
(2) You only need to determine whether the given constraint is related to length, without considering other constraints of the user instruction.
(3) Output your analysis and conclusion strictly following the format below:
Analysis: …(Provide a detailed analysis to determine whether the given constraint is related to length)
Conclusion: …
1. (If the constraint is NOT related to length)
”””json
   ”Length Constraint” : False
2. (If the constraint is related to length)
”””json
   ”Length Constraint” : True,
   ”Extracted Segments” : [
       ”Length Requirement within the Constraint”: …(Provide the length requirement within the constraint. Be sure to extract the original text from the constraint without making any modification),
       ”Corresponding Segment in Response”: …(Provide the segment of the response that should follow this requirement without making any modification. If no corresponding segment exists in the response, output ”No corresponding segment exists.”)
       ”Length Requirement within the Constraint”: …,
       ”Corresponding Segment in Response”: …
     … (Analogously, extracting all segments in the response that should follow the length requirements of the given constraint)
Table 11: The prompt template for length constraint identification and segment extraction.
You are a fair judge, specializing in evaluating the quality of an AI assistant’s responses to user instructions.
We will provide you with a user instruction, a specific constraint from the user instruction, and the AI assistant’s response. Please think carefully and provide a detailed analysis of whether the AI assistant’s response follows the given constraint.
The given constraint is related to length. We will also provide a JSON data, which contains multiple items. Each item consists of three elements: (1) a length requirement within the constraint, (2) a segment of the responses that should follow the length requirement, and (3) the actual length of the segment. It is important that you consider the ”actual length provided in the JSON data” as the golden-truth length, which is absolutely accurate. Do not make your own calculation about “the length of the segment”. You must rely solely on the ”actual length” as the basis for length validation. Output your explanation and judgment for the given constraint based on the actual length provided in the JSON data, strictly following the format below:
[The Start of Constraint]
Constraint: … (Directly output the first constraint from the checklist here with no modifications)
Explanation: … (Provide a thorough, step-by-step analysis of whether the AI assistant’s response follows this constraint, referencing specific details of the response)
Judgment: [[The AI assistant’s response follows this constraint]] or [[The AI assistant’s response does not follow this constraint]]
[The End of Constraint]
Remember the following points:
(1) You should only focus on whether the response follows the constraint, not on the overall quality of the response. Other shortcomings in the response (such as not fulfilling the constraint to a high standard) should not affect your judgment unless the constraint is not followed.
(2) Your judgment should be as strict as possible. You should output ”[[The AI assistant’s response follows this constraint]]” only if the response completely follows every part of the constraint as stated, without any omission or mistake.
(3) Independently analyzing the given constraint, without taking into account whether other constraints are followed.
(4) The given constraint may contain both length requirements and other requirements. For length requirements, you must rely solely on the ”actual length provided in the JSON data” as the basis for length validation. For other requirements, please refer to the above principles and carefully evaluate them by yourself.
(5) When generating your explanation, it is forbidden to mention phrases such as ”according to the given JSON data” or ”based on the provided actual length”. Your explanation should convey that the length is calculated independently by you, without referencing the given JSON data.
Here are the user instruction, the constraint, the AI assistant’s response, and the JSON data about length:
[The Start of User Instruction]
{instruction}
[The End of User Instruction]
[The Start of the Constraint]
{constraint}
[The End of the Constraint]
[The Start of AI Assistant’s Response]
{response}
[The End of AI Assistant’s Response]
[The Start of The Json Data]
{json_data}
[The End of The Json Data]
Table 12: The prompt template for rule-augmented critique revision.
Below, an instruction, a corresponding model response, and a constraint checklist for the instruction will be provided. The checklist contains some of the constraints within the instruction. Your task is to verify whether the model response follows each constraint in the checklist, choose ”Followed” or ”Not Followed”, and provide a brief justification.
Task Details:
1. Please analyze whether the response follows each constraint listed in the given checklist, providing a judgment for each constraint respectively.
2. Your judgments must be strict. Only responses that fully satisfy a constraint can be judged as ”Followed”. If there is any omission or error regarding a constraint, it must be judged as ”Not Followed”.
3. Please focus exclusively on the constraints within the given checklist. It is unnecessary to consider whether the response follows any other constraints beyond the checklist.
4. When judging the following of each constraint, your judgement should consider the complete context of the instructions, rather than interpreting the constraint in isolation.
{in-context examples}
[Instruction]
{instruction}
[Model Response]
{model_response}
[Constraint Checklist]
{checklist}
Your choice for the first constraint in the checklist: {option}  A. Followed B. Not Followed
Your justification: {justification}
Your choice for the second constraint in the checklist: {option}  A. Followed B. Not Followed
Your justification: {justification}
Table 13: Human annotation guideline for constraint verification. The red part is the information provided to the annotators, and the blue part is content that requires the annotators to make annotations.

### v2

@@
 Abstract
-Instruction following is a fundamental ability of Large Language Models (LLMs), requiring their generated outputs to follow multiple constraints imposed in input instructions.
+Instruction-following is a fundamental ability of Large Language Models (LLMs), requiring their generated outputs to follow multiple constraints imposed in input instructions.
 Numerous studies have attempted to enhance this ability through preference optimization or reinforcement learning based on reward signals from LLM-as-a-Judge.
-However, existing evaluation models for instruction following still possess many deficiencies, such as substantial costs and unreliable assessments.
-To this end, we propose IF-Critic, an LLM critic that can provide efficient and reliable assessments of constraint following in the instructions.
+However, existing evaluation models for instruction-following still possess many deficiencies, such as substantial costs and unreliable assessments.
+To this end, we propose IF-Critic, an LLM critic for fine-grained, efficient, and reliable instruction-following evaluation.
 We first develop a checklist generator to decompose instructions and generate constraint checklists.
 With the assistance of the checklists, we collect high-quality critique training data through a multi-stage critique filtering mechanism and employ a constraint-level preference optimization method to train IF-Critic.
-Extensive experiments demonstrate that the evaluation performance of IF-Critic can beat strong LLM-as-a-Judge baselines, including Deepseek-R1 and o4-mini.
-With the scalable reward signals provided by IF-Critic, LLMs can achieve substantial performance gains in instruction-following optimization under lower computational overhead compared to strong LLM critic baselines.
-Work done when this author interned at Zhipu AI.Equal contribution.Corresponding author.
-Codehttps://github.com/thu-coai/IF-CRITIC
-Introduction
-Large language models (LLMs) have demonstrated remarkable capabilities in various NLP tasks (Zhao et al.
-2023).
-Among these, instruction following stands out as one of the most crucial requirements for LLM applications (Ouyang et al.
-2022).
-In the real-world use of LLMs, nearly all tasks are formulated as instruction following, where human instructions specify task requirements and impose corresponding constraints on the model output (Jiang et al.
-2024).
-Accurately following instructions serves as a key factor in ensuring the helpfulness and reliability of LLMs (Huang et al.
-2024).
-However, LLMs still have not excelled in instruction following, especially when dealing with complex instructions that involve numerous constraints (Jiang et al.
-2024; Wen et al.
-2024).
+Extensive experiments show that the evaluation performance of IF-Critic can beat strong LLM-as-a-Judge baselines, including o4-mini and Gemini-3-Pro.
+With the reward signals provided by IF-Critic, LLMs can achieve substantial performance gains in instruction-following optimization under lower computational overhead compared to strong LLM critic baselines.
+Our code and model are available at https://github.com/thu-coai/IF-CRITIC.
+IF-Critic: Towards a Fine-Grained LLM Critic for Instruction-Following Evaluation
+Bosi Wen1,†,∗   Yilin Niu2,∗  Cunxiang Wang2  Pei Ke3  Xiaoying Ling2 Ying Zhang2  Aohan Zeng4  Hongning Wang1   Minlie Huang1,‡ 1The Conversational Artificial Intelligence (CoAI) Group, Tsinghua University 2Zhipu AI  3University of Electronic Science and Technology of China 4The Knowledge Engineering Group (KEG), Tsinghua University wbs23@mails.tsinghua.edu.cn, aihuang@tsinghua.edu.cn
+Work done when this author interned at Zhipu AI.Equal contributionCorresponding author
+1 Introduction
+Large language models (LLMs) have demonstrated remarkable capabilities in various NLP tasks Zhao et al.
+(2023).
+Among these, instruction-following stands out as one of the most crucial requirements for LLM applications Ouyang et al.
+(2022).
+In the real-world use of LLMs, nearly all tasks are formulated as instruction-following, where human instructions specify task requirements and impose corresponding constraints on the model output Jiang et al.
+(2024).
+Accurately following instructions serves as a key factor in ensuring the helpfulness and reliability of LLMs Huang et al.
+(2024).
+However, LLMs still have not excelled in instruction-following, especially when dealing with complex instructions that involve numerous constraints Jiang et al.
+(2024).
 Numerous research attempts have been made to improve the instruction-following ability of LLMs.
-They often leverage LLMs to synthesize complex instructions and perform preference optimization or reinforcement learning based on the reward signals from LLM-as-a-Judge (Sun et al.
-2024; Zhang et al.
-2024c; He et al.
-2024a; Ren et al.
-2025; Liu et al.
-2025b; Peng et al.
-2025).
+They often leverage LLMs to synthesize complex instructions and perform preference optimization or reinforcement learning based on the reward signals from LLM-as-a-Judge Zhang et al.
+(2025); He et al.
+(2024a); Ren et al.
+(2025); Liu et al.
+(2025b); Peng et al.
+(2025).
 Figure 1: A usage example of IF-Critic: Given an instruction and a response, a checklist generator first decomposes the instruction to generate a constraint checklist.
-Then, IF-Critic can provide fine-grained evaluations for the response with respect to its following of all included constraints.
-Although these works have achieved promising results, the usage of LLM-as-a-Judge for instruction-following still faces two major yet underexplored challenges: (1) Substantial Costs: Existing methods typically rely on proprietary LLMs (e.g., GPT-4o) or large reasoning models (e.g., QwQ-32B) to provide evaluation results as reward signals, which is costly and difficult to scale up in training data construction.
-(2) Unreliable Assessments: Current LLMs still struggle to accurately assess instruction-following (Zhang et al.
-2024b), exhibiting low recall in error detection (Kamoi et al.
-2024) and limited counting ability (Zhang, Cao, and You 2024; Fu et al.
-2024).
+Then, IF-Critic can provide fine-grained evaluations for the response with respect to its following of all included constraints in one inference pass.
+Despite these efforts, the usage of LLM-as-a-Judge for instruction-following still faces two major yet underexplored challenges: (1) Substantial Costs: Existing methods typically rely on proprietary LLMs (e.g., GPT-4o) or large reasoning models (e.g., QwQ-32B) to provide evaluation results as reward signals, which is costly and difficult to scale up in model training.
+(2) Unreliable Assessments: Current LLMs still struggle to accurately assess instruction-following Zhang et al.
+(2024c), exhibiting low recall in error detection Kamoi et al.
+(2024) and limited counting ability Zhang et al.
+(2024b); Fu et al.
+(2024).
 These limitations introduce noise to the evaluation and restrict the gains in model optimization.
-To tackle the above challenges, some works attempt to introduce handcrafted constraints that can be automatically verified by code execution in data construction, aiming to enhance the efficiency and reliability of evaluation (Dong et al.
-2025; Kim et al.
-2025; Liu et al.
-2025b; Peng et al.
-2025).
-However, these constraints are often homogeneous and atomized, failing to cover the diversity and complexity of human instructions, such as constraints composition (Wen et al.
-2024).
-In this work, we propose IF-Critic, an LLM critic with 14B parameters tailored for instruction-following evaluation.
+To tackle these challenges, some works attempt to introduce handcrafted code-verifiable constraints in data construction, aiming to enhance evaluation efficiency and reliability Dong et al.
+(2025); Kim et al.
+(2025); Li et al.
+(2025); Peng et al.
+(2025).
+However, these constraints are often homogeneous and atomized, failing to cover the diversity and complexity of human instructions, such as constraint composition Wen et al.
+(2024).
+In this work, we propose IF-Critic, an LLM critic tailored for instruction-following evaluation.
 To enhance inference efficiency and obtain a more holistic and granular perception of the instructions, IF-Critic adopts a novel checklist-guided critique generation paradigm.
@@
 Then, IF-Critic can evaluate the following of all included constraints in one inference pass, eliminating the need to generate separate judgments for different constraints multiple times.
-To train IF-Critic, we elaborately prompt Deepseek-R1 (Guo et al.
-2025a) to generate expert critiques guided by the checklists, then employ a multi-stage critique filtering mechanism to select the highest-quality critique for each constraint, which mitigates noise and significantly enhances the reliability of critiques for various constraints, including those involving counting.
-After fine-tuning on this data, we leverage the property that the critiques contain multi-constraint evaluation results in the checklist, further improving IF-Critic via a novel constraint-level preference optimization method.
-This method performs fine-grained, constraint-level comparisons between positive and negative critiques using Direct Preference Optimization (DPO) (Rafailov et al.
-2023), reinforcing the perception of crucial preference information and leading to better alignment between IF-Critic and the expert critiques.
-Empirically, IF-Critic consistently outperforms various LLM-as-a-Judge baselines on four instruction-following meta-evaluation benchmarks, including the latest state-of-the-art LLMs Deepseek-R1 and o4-mini (Jaech et al.
-2024).
-Leveraging the fine-grained evaluation results from IF-Critic as reward signals, we adapt DPO and Group Relative Policy Optimization (GRPO) (Shao et al.
-2024) to improve the instruction-following ability of LLMs.
+To train IF-Critic, we elaborately prompt Deepseek-R1 Guo et al.
+(2025a) to generate expert critiques guided by the checklists, then employ a multi-stage critique filtering mechanism to select the highest-quality critique for each constraint, which significantly mitigates noise and enhances the reliability of critiques, including those involving counting.
+After fine-tuning on this data, we leverage the property that the critiques contain multi-constraint evaluation results, further improving IF-Critic via a novel constraint-level preference optimization method.
+This method performs fine-grained, constraint-level comparisons between positive and negative critiques using Direct Preference Optimization (DPO) Rafailov et al.
+(2023), reinforcing the perception of crucial preference information and leading to better alignment between IF-Critic and the expert critiques.
+Empirically, IF-Critic consistently outperforms various LLM-as-a-Judge baselines on four instruction-following meta-evaluation benchmarks, including the latest state-of-the-art LLMs o4-mini Jaech et al.
+(2024) and Gemini-3-Pro Team et al.
+(2023).
+Leveraging the fine-grained evaluation results from IF-Critic as reward signals, we adapt DPO and Group Relative Policy Optimization (GRPO) Shao et al.
+(2024) to improve the instruction-following ability of LLMs.
 Experiments demonstrate that IF-Critic brings substantially greater improvements over strong baselines with significantly less computational overhead, highlighting its effectiveness and practicability.
@@
   • We propose a novel framework for the development of instruction-following critics, including the checklist-guided critique generation paradigm, the multi-stage critique filtering mechanism for high-quality training data collection, and the constraint-level preference optimization method for model training.
-  • We introduce IF-Critic, the first instruction-following critic with 14B parameters, developed through our framework.
-Experiments show that IF-Critic achieves superior evaluation performance compared to strong LLM-as-a-Judge baselines, including Deepseek-R1 and o4-mini.
-  • We validate the potential of IF-Critic to provide scalable reward signals and improve the instruction-following abilities of LLMs using DPO and GRPO.
+  • We introduce IF-Critic, the first instruction-following critic developed through our framework.
+Experiments show that IF-Critic achieves superior evaluation performance compared to strong LLM-as-a-Judge baselines, including o4-mini and Gemini-3-Pro.
+  • We validate that IF-Critic can provide scalable reward signals to improve the instruction-following abilities of LLMs using DPO and GRPO.
 IF-Critic can bring substantially greater improvements under lower computational overhead compared to strong baselines.
-Related Work
+2 Related Work
 LLM-based Evaluation.
-Utilizing LLMs as text evaluators has gradually become prevalent (Wang et al.
-2023a; Chen et al.
-2023; Ke et al.
-2023) due to their advantages in flexibility, interpretability, and generalization.
-Although proprietary LLMs have demonstrated excellent evaluation capabilities (Zheng et al.
-2023; Wang et al.
-2023a), their practical application is limited by issues such as high cost and potential data leakage (Ke et al.
-2024).
-To reduce the reliance on proprietary LLMs, a series of works attempt to train smaller LLMs specialized in evaluation using the data generated from humans or proprietary LLMs (Kim et al.
-2024a; Li et al.
-2024; Ke et al.
-2024; Hu et al.
-2024; Kim et al.
-2024b).
+Utilizing LLMs as text evaluators has gradually become prevalent Wang et al.
+(2023a); Chen et al.
+(2023); Ke et al.
+(2023) due to their advantages in flexibility, interpretability, and generalization.
+Although proprietary LLMs have demonstrated excellent evaluation capabilities Zheng et al.
+(2023); Wang et al.
+(2023a), their practical application is limited by issues such as high cost and potential data leakage Ke et al.
+(2024).
+To reduce the reliance on proprietary LLMs, a series of works attempt to train smaller LLMs specialized in evaluation using the data generated from humans or proprietary LLMs Kim et al.
+(2024a); Li et al.
+(2024); Ke et al.
+(2024); Hu et al.
+(2024); Kim et al.
+(2024b).
 Distinct from these works that focus on evaluating overall text quality, we develop an evaluation model that provides constraint-level, fine-grained assessments of instruction-following and leverages them as reward signals to enhance the instruction-following ability of LLMs.
-Instruction Following Optimization.
-Instruction following remains a critical factor in determining the practicality of LLMs (Liu et al.
-2023; Lou, Zhang, and Yin 2024), driving extensive studies to enhance this ability.
-Earlier works primarily focus on following simple instructions that mostly consist of a single constraint (Chiang et al.
-2023; Wang et al.
-2023c).
-With the expanding application of LLMs, recent studies place greater emphasis on complex instruction-following that involves numerous constraints.
-They typically leverage LLMs to synthesize complex instructions and perform preference optimization or reinforcement learning based on the reward signals from LLM-as-a-Judge (Sun et al.
-2024; Zhang et al.
-2024c; Cheng et al.
-2025; Ren et al.
-2025; Liu et al.
-2025b; Peng et al.
-2025).
-Specifically, SPaR (Cheng et al.
-2025) employs self-play tree search refinement through a fine-tuned refiner to construct DPO training data.
-RECAST (Liu et al.
-2025b) categorizes constraint synthesis into soft and hard constraints, leveraging GPT-4o and code execution, respectively, to obtain evaluation results for GRPO training.
+instruction-following Optimization.
+As LLMs have been increasingly applied to real-world complex tasks, following complex instructions that involve numerous constraints emerges as a key determinant of their practical utility Liu et al.
+(2023); Lou et al.
+(2024), driving extensive studies to enhance this ability of LLMs.
+They typically leverage LLMs to synthesize complex instructions and perform preference optimization or reinforcement learning based on the reward signals from LLM-as-a-Judge Sun et al.
+(2024); Zhang et al.
+(2025); Cheng et al.
+(2025); Ren et al.
+(2025); Liu et al.
+(2025b); Peng et al.
+(2025).
+Specifically, SPaR Cheng et al.
+(2025) employs self-play tree search refinement through a fine-tuned refiner to construct DPO training data.
+RECAST Liu et al.
+(2025b) categorizes soft and hard constraints, leveraging GPT-4o and code execution, respectively, to obtain evaluation results for GRPO training.
 However, the effectiveness of existing methods is restricted by the high costs and limited instruction-following evaluation ability of LLM-as-a-Judge, and code-verifiable constraints fail to cover the diversity and complexity of human instructions.
-IF-Critic Framework
+3 Methodology
 Figure 2: The pipeline of IF-Critic development. The left section illustrates the process of critique training data construction, while the right section presents the process of training IF-Critic.
-Task Definition and Method Overview
-Distinct from previous LLM critics that typically generate evaluation results for a single criterion or question (Wang et al.
-2024a; Vu et al.
-2024), IF-Critic leverages a constraint checklist to guide critique generation and evaluates the following of all constraints included in the checklist within one inference pass.
-Specifically, given an instruction $x$, corresponding response $y$, and a constraint checklist $\{c_{k}\}^{n}_{k=1}$ which contains all constraints within $x$, IF-Critic can provide a critique $C=\bigcup_{k=1}^{n}(e_{k},j_{k})$ for $y$, where each segment $(e_{k},j_{k})$ consists of an explanation $e_{k}$ and a binary judgment $j_{k}$ (0 or 1) for the respective constraint following.
+3.1 Task Definition and Method Overview
+Distinct from previous LLM critics that typically generate evaluation results for a single criterion or question Vu et al.
+(2024); Wang et al.
+(2025), IF-Critic leverages a constraint checklist to guide critique generation and evaluates the following of all included constraints in one inference pass.
+Specifically, given an instruction $x$, corresponding response $y$, and a constraint checklist $\{c_{k}\}^{n}_{k=1}$ which contains all constraints within $x$, IF-Critic can provide a critique $C=\bigcup_{k=1}^{n}(e_{k},j_{k})$ for $y$, where each segment $(e_{k},j_{k})$ consists of an explanation $e_{k}$ and a binary judgment $j_{k}$ (0 or 1) for the following of the respective constraint.
 This paradigm not only enhances inference efficiency but also equips IF-Critic with a more holistic and granular perception of the instructions, thus improving the reliability of instruction-following assessments.
@@
 IF-Critic can be leveraged to provide reliable reward signals, improving the instruction-following ability of LLMs via DPO or GRPO, as validated in our experiments.
-Instruction and Response Collection
-To obtain diverse instructions for critic training, we collect instructions from real application scenarios and utilize LLMs to categorize them according to the task taxonomy of CritiqueLLM (Ke et al.
-2024), which encompasses 10 categories and covers diverse NLP applications in real-world scenarios.
+3.2 Instruction and Response Collection
+To obtain diverse instructions for critic training, we collect instructions from real-world application scenarios and utilize LLMs to categorize them according to the task taxonomy of CritiqueLLM Ke et al.
+(2024), which encompasses 10 categories and covers diverse NLP applications in real-world scenarios.
 To improve data quality, we prompt LLMs to score instruction quality and develop a classifier to assess their constraint complexity, which is trained with a small amount of human annotation data.
 Finally, we balance the amount of instruction across task categories and acquire 55K high-quality instructions with relatively high complexity.
-Then, we collect LLM responses from 15 representative models, which possess varying abilities in instruction-following, including GPT-4o (OpenAI 2023), GPT-4o-mini, seven versions of GLM-4 (GLM et al.
-2024), three versions of Qwen2.5 (Yang et al.
-2024), Deepseek-R1 (Guo et al.
-2025a), Doubao1.5 (Doubao Team 2025), and Claude3.7-sonnet (Anthropic 2023).
+Then, we collect LLM responses from 15 representative models, which possess varying abilities in instruction-following.
+The model list is provided in Appendix A.
 For each instruction, we randomly select two models to generate responses, resulting in a total of 110k evaluation inputs, each comprising an instruction and its corresponding LLM response.
-Critique Training Data Construction
-To assist IF-Critic in providing fine-grained assessments of instruction-following, we first develop a checklist generator to decompose the constraints within instruction $x$ and generate a constraint checklist $\{c_{k}\}^{n}_{k=1}$.
-We leverage a powerful LLM, Deepseek-R1111https://huggingface.co/deepseek-ai/DeepSeek-R1, for automatic checklist annotation in our collected instructions and utilize its generated checklists to fine-tune the checklist generator.
-Manual inspection of 1,000 test samples shows the accuracy of our checklist generator exceeds 99%, indicating its reliability.
+3.3 Critique Training Data Construction
+To assist IF-Critic in fine-grained evaluation, we first develop a checklist generator to decompose the constraints within instruction $x$ and generate a constraint checklist $\{c_{k}\}^{n}_{k=1}$.
+We leverage a powerful LLM, Deepseek-R1  [^1] https://huggingface.co/deepseek-ai/DeepSeek-R1, for automatic checklist annotation in our collected instructions.
+Then, its generated checklists are utilized to fine-tune the generator, which supports efficient checklist generation after deployment.
+Manual inspection of 1,000 test samples for the checklist generator shows that 99.29% of generated constraints and 97.50% of entire checklists are correct, indicating its reliability.
 With the assistance of the checklists, we proceed to employ Deepseek-R1 for expert critique annotation via a carefully designed prompt, which elicits a concise and specific explanation before the judgment for each constraint.
-Detailed prompts are in the Appendix.
-For each response, we collect $N$ critiques and adapt a multi-stage critique filtering mechanism to select the highest-quality critique for each constraint in the checklist, which is then used for training IF-Critic.
+Detailed prompts are in the Appendix B.
+For each response, we collect $N$ expert critiques and adapt a multi-stage critique filtering mechanism to select the highest-quality critique for each constraint in the checklist, which is then used for training IF-Critic.
 The mechanism comprises four stages.
-Specifically, we first employ cross-model and rule-augmented verification to filter out noise in critiques caused by two major limitations of LLM-as-a-Judge: potential evaluation bias and limited counting ability, respectively.
-Subsequently, we select the highest-quality judgements and explanations from the resulting noise-free critiques as the final critiques based on consistency, significantly enhancing their reliability.
+We first employ cross-model and rule-augmented verification to filter out noise in critiques caused by two major limitations of LLM-as-a-Judge: potential evaluation bias and limited counting ability, respectively.
+Subsequently, we select the highest-quality judgements and explanations from the resulting noise-free critiques as the final critiques based on consistency, further enhancing their reliability.
 Cross-Model Verification.
-Previous works point out that LLM-as-a-Judge may suffer from potential biases (Zheng et al.
-2023; Ye et al.
-2025a), undermining the reliability of its generated critiques.
+Previous works point out that LLM-as-a-Judge may suffer from potential biases Zheng et al.
+(2023); Ye et al.
+(2025a), undermining the reliability of its generated critiques.
 To mitigate this, we utilize two additional LLMs, GLM-4-Plus and Qwen2.5-72B-Instruct, to independently verify two aspects of the expert critiques for each constraint: (1) The correctness of the explanation, and (2) The consistency between the explanation and judgment.
-Detailed prompts can be found in the Appendix.
+Detailed prompts can be found in Appendix B.
 Any critiques that fail verification by any model in any aspect are strictly filtered, accounting for 11.3% of our data.
 Rule-Augmented Verification.
-Considering that LLMs often struggle with accurate counting (Zhang, Cao, and You 2024; Fu et al.
-2024), we introduce rule-augmented verification to enhance the reliability of critiques involving length constraints (e.g., each title must be 8 characters long).
-Methodologically, we first prompt Qwen2.5-72B-Instruct to identify length constraints and extract the segments from the responses that are relevant to these constraints.
-Subsequently, we employ Python scripts to automatically identify length information within these segments, and require Deepseek-R1 to revise its critiques with the assistance of this information.
-The prompts for these steps are in the Appendix.
+Considering that LLMs often struggle with accurate counting Zhang et al.
+(2024b); Fu et al.
+(2024), we introduce rule-augmented verification to enhance the reliability of critiques involving length constraints (e.g., each title must be 8 characters long).
+Methodologically, we first prompt Qwen2.5-72B-Instruct to identify length constraints and extract relevant segments from the responses.
+Subsequently, we employ Python scripts to identify length information within these segments automatically.
+This information then guides Deepseek-R1 to revise its critiques.
+The prompts for these steps are in the Appendix B.
 This approach integrates the flexibility of LLM-based evaluation with the precision of rule-based evaluation, thereby effectively handling the diverse length constraints in human instructions.
 Final Judgement Selection.
-The correctness of judgment is vitally important in instruction-following evaluation.
-Given that critique generation is analogous to a Chain-of-Thought (CoT) (Wei et al.
-2022) reasoning process, we incorporate self-consistency (Wang et al.
-2023b), a widely used decoding strategy in CoT reasoning, to select final judgments and improve their correctness.
+Given that critique generation is analogous to a Chain-of-Thought (CoT) Wei et al.
+(2022) reasoning process, we incorporate self-consistency Wang et al.
+(2023b), a widely used decoding strategy in CoT reasoning, to select final judgments and improve their correctness.
 Specifically, for each constraint $c_{k}$, we select the final judgment $j_{k}^{*}$ via majority voting among multiple expert critiques.
-Furthermore, previous works demonstrate that confidence, which is calculated as the proportion of the majority result to the total number of samplings, can serve as an important indicator of answer reliability (Xiong et al.
-2024; Hu et al.
-2024; Prasad et al.
-2024), with higher confidence generally implying greater reliability.
+Furthermore, previous works show that confidence, which is calculated as the proportion of the majority result to the total samples, is an important indicator of answer reliability Xiong et al.
+(2024); Hu et al.
+(2024); Prasad et al.
+(2024), with higher confidence generally implying greater reliability.
 Inspired by this, we discard the final critiques of constraints with confidence lower than 0.75 in our experiment, which empirically balances data retention and reliability.
@@
 After obtaining the correct final judgments, the last step is to select the highest-quality explanations that support these judgments.
-To achieve this, we devise a selection strategy inspired by Minimum Bayes Risk Decoding (Bickel and Doksum 2015).
-Specifically, for each constraint $c_{k}$, we first construct a hypothesis set $\mathcal{H}_{k}$ containing all explanations whose corresponding judgments align with the final judgments $j_{k}^{*}$.
+To achieve this, we devise a selection strategy inspired by Minimum Bayes Risk Decoding Bickel and Doksum (2015): For each constraint $c_{k}$, we first construct a hypothesis set $\mathcal{H}_{k}$ containing all explanations whose corresponding judgments align with the final judgments $j_{k}^{*}$.
 Then, we utilize a text similarity measure $u(\cdot)$ to calculate the consistency between each pair of explanations in $\mathcal{H}_{k}$.
@@
 $e_{k}^{*}=\arg\max_{e\in\mathcal{H}_{k}}\frac{1}{|\mathcal{H}_{k}|}\sum_{\widetilde{e}\in\mathcal{H}_{k}}u(\widetilde{e},e)$ (2)
-IF-Critic Training
+Through these rigorous filtering processes, we manually inspect the quality of 70 final critiques with 353 constraints and find that 96.03% judgments and 92.35% explanations are completely correct, demonstrating their high quality.
+3.4 IF-Critic Training
 After constructing the final critique training dataset $D=\{C_{i}=\bigcup_{k=1}^{n_{i}}(e_{ik}^{*},j_{ik}^{*})\}_{i=1}^{|D|}$, we partition $D$ into two subsets $D_{\text{sft}}\cup D_{\text{ref}}$ and employ a two-stage paradigm to train IF-Critic, comprising supervised fine-tuning (SFT) and constraint-level preference optimization.
-In the first stage, we utilize a critique generation prompt template $p$ to concatenate the instruction $x$, response $y$, and checklist $\{c_{k}\}^{n}_{k=1}$, applying a standard SFT loss on vanilla LLM $\theta$:
+In the first stage, we utilize a critique generation prompt $p$ to concatenate the instruction $x$, response $y$, and checklist $\{c_{k}\}^{n}_{k=1}$, applying a standard SFT loss on vanilla LLM $\theta$:
 $\mathcal{L}_{\text{SFT}}=-\sum_{i=1}^{|D_{\text{sft}}|}\text{log}P_{\theta}(C_{i}|p_{i})$ (3)
-After obtaining the SFT critic, we seek to further enhance its capability via preference optimization.
-Existing methods usually sample multiple independent responses from the policy model and utilize response-level rewards to select chosen and rejected examples (Yuan et al.
-2024; Ye et al.
-2025b; Yu et al.
-2025).
-However, since the critiques in our task involve multiple segments for different constraints in the checklist, only the segments where chosen and rejected critiques differ in judgments encapsulate crucial preference information.
-Other segments may introduce unnecessary interference and dilute important supervision signals, leading to insufficient optimization (Chen et al.
-2024; Cheng et al.
-2025).
+After obtaining the SFT critic, we further enhance its ability via preference optimization.
+Existing methods usually sample multiple independent responses from the policy model and utilize response-level rewards to select chosen and rejected examples Yuan et al.
+(2024); Ye et al.
+(2025b); Yu et al.
+(2025).
+However, since the critiques in our task involve multiple segments for different constraints, only the segments where chosen and rejected critiques differ in judgments encapsulate crucial preference information.
+Other segments may introduce unnecessary interference and dilute important supervision signals, leading to insufficient optimization Chen et al.
+(2024); Cheng et al.
+(2025).
 To mitigate this, we introduce fine-grained, constraint-level comparisons in preference pairs construction, thus reinforcing the perception of crucial preference information.
-Specifically, for each instruction $x$ in $D_{\text{ref}}$, we first sample $M$ critiques from the SFT critic.
+Specifically, for each evaluation input $p$ in $D_{\text{ref}}$, we first sample $M$ critiques from the SFT critic.
 Then, we identify the best self-sampled explanation $\hat{e}_{k}$ for each constraint $c_{k}$ whose judgements align with the expert final judgement $j^{*}_{k}$ in a manner analogous to Equation 1 and 2 (if any).
 Finally, self-sampled critiques with at least one judgment that misaligns with the expert final critique are selected as rejected critiques $C_{l}=\bigcup_{k=1}^{n}(e_{k},j_{k})$.
-Retaining the segments that align with the expert final critique while replacing other misaligned segments with the best self-sampled explanations $\hat{e}_{k}$ to construct chosen critique $C_{w}$:
+Retaining the segments that align with the expert final critique while replacing other misaligned segments with the best self-sampled explanations $\hat{e}_{k}$ and expert final judgement $j^{*}_{k}$ to construct chosen critiques $C_{w}$:
 $(e^k,jk),jkjk\displaystyle C_{w}=\bigcup_{k=1}^{n}$ (4)
@@
 Consequently, we construct a preference dataset $D_{\text{dpo}}$ where the preference pairs only differ in the segments with inconsistent judgments, thereby reinforcing crucial preference information.
-Moreover, self-sampled critiques are closer to the SFT critic decoding space than the expert critique, which is generally beneficial for optimization (Guo et al.
-2024).
+Moreover, self-sampled critiques are closer to the SFT critic decoding space than the expert critique, which is generally beneficial for optimization Guo et al.
+(2024).
 Finally, we conduct preference optimization via the following DPO loss:
@@
 where both $\pi_{\theta}$ (i.e., the trained policy model) and $\pi_{\text{ref}}$ (i.e., the fixed reference model) are initialized from the SFT critic.
-Evaluation Results Utilization
+3.5 Evaluation Results Utilization
 Leveraging the evaluation results of IF-Critic as reward signals, we can apply DPO or GRPO algorithms to improve the instruction-following abilities of LLMs.
@@
 Then, we can use this reward to construct DPO training data or to conduct GRPO training.
-Experiments
-Instruction-Following Evaluation
+4 Experiments
+4.1 Instruction-Following Evaluation
 Evaluation Benchmarks.
-Considering the scarcity of existing instruction-following meta-evaluation datasets, we sample instructions from four benchmarks: EvalCritic, CFBench (Zhang et al.
-2024a), TRACE (Zhang et al.
-2024c), and Multi-IF (He et al.
-2024b) to construct meta-evaluation datasets.
-EvalCritic is a random split of the dataset we used for instruction-following optimization [^2] We elaborate on this dataset in the following ”Instruction-Following Optimization” section., whose distribution of instructions is distinct from the training data of IF-Critic.
+Given the scarcity of existing instruction-following meta-evaluation datasets, we sample instructions from four benchmarks: EvalCritic, CFBench Zhang et al.
+(2024a), TRACE Zhang et al.
+(2025), and Multi-IF He et al.
+(2024b) to construct meta-evaluation datasets.
+EvalCritic is a random split of the dataset we used for instruction-following optimization (§4.2) , whose distribution of instructions differs from IF-Critic training data.
 For the first three benchmarks, we randomly sample 150 instructions.
 For Multi-IF, we select the first round from the Chinese version, comprising 454 instructions.
-Responses for each instruction are generated by 2 out of 6 LLMs with varying capabilities.
+Responses for each instruction are generated by 2 of 6 LLMs with varying capabilities.
 We use our checklist generator to generate constraint checklists for EvalCritic and directly adapt the original checklists for other benchmarks.
-For the first three benchmarks, two annotators are required to independently judge the following of each constraint, with a third inspector conducting cross-validation.
+For the first three benchmarks, two annotators independently judge the following of each constraint, with a third inspector conducting cross-validation.
 Any discrepancies will be discussed to reach a consensus.
 For Multi-IF, we use the provided verification code for evaluation.
-Across these benchmarks, there are 1411 / 1070 / 1634 / 1220 constraints, respectively, of which 406 / 252 / 297 / 320 are labeled as ”not followed”.
-Prompt Model EvalCritic CFBench TRACE Multi-IF (zh) Avg.
+Across these benchmarks, there are 1411 / 1070 / 1634 / 1220 constraints, respectively, of which 406 / 252 / 297 / 320 are labeled as "not followed".
+Prompt Model EvalCritic CFBench TRACE Multi-IF Avg.
 Positive Negative Positive Negative Positive Negative Positive Negative Average
 F1 F1 F1 F1 F1 F1 F1 F1 F1
-Constraint-Level o4-mini 0.887 0.717 0.927 0.755 0.923 0.599 0.960 0.890 0.832
+Constraint-Level Gemini-3-Pro 0.887 0.744 0.934 0.794 0.932 0.704 0.962 0.898 0.857
+o4-mini 0.887 0.717 0.927 0.755 0.923 0.599 0.960 0.890 0.832
+GPT-4.1 0.883 0.608 0.919 0.689 0.924 0.597 0.932 0.801 0.794
 Deepseek-R1 0.863 0.686 0.905 0.718 0.883 0.555 0.933 0.824 0.796
-GPT-4.1 0.883 0.608 0.919 0.689 0.924 0.597 0.932 0.801 0.794
-GLM-4-Plus 0.873 0.526 0.900 0.563 0.926 0.518 0.913 0.682 0.738
-Qwen2.5-72B-Instruct 0.854 0.362 0.891 0.459 0.915 0.374 0.896 0.590 0.668
+QwQ-32B 0.856 0.675 0.900 0.684 0.901 0.595 0.931 0.795 0.792
+Qwen-2.5-72B-Instruct 0.854 0.362 0.891 0.459 0.915 0.374 0.896 0.590 0.668
 Llama-3.3-70B-Instruct 0.869 0.486 0.904 0.612 0.916 0.426 0.921 0.756 0.736
-QwQ-32B 0.856 0.675 0.900 0.684 0.901 0.595 0.931 0.795 0.792
 GLM-4-32B-0414 0.867 0.451 0.900 0.531 0.913 0.384 0.898 0.610 0.694
-Checklist-Level o4-mini 0.911 0.753 0.935 0.761 0.925 0.639 0.964 0.900 0.849
+Checklist-Level Gemini-3-Pro 0.902 0.742 0.945 0.808 0.932 0.655 0.960 0.892 0.855
+o4-mini 0.911 0.753 0.935 0.761 0.925 0.639 0.964 0.900 0.849
+GPT-4.1 0.883 0.560 0.918 0.637 0.923 0.517 0.933 0.799 0.771
 Deepseek-R1 0.895 0.716 0.928 0.725 0.910 0.579 0.938 0.828 0.815
-GPT-4.1 0.883 0.560 0.918 0.637 0.923 0.517 0.933 0.799 0.771
-GLM-4-Plus 0.856 0.354 0.903 0.524 0.916 0.369 0.906 0.649 0.685
-Qwen2.5-72B-Instruct 0.849 0.278 0.894 0.424 0.914 0.330 0.889 0.548 0.641
+QwQ-32B 0.892 0.663 0.922 0.716 0.922 0.570 0.933 0.793 0.801
+Qwen-2.5-72B-Instruct 0.849 0.278 0.894 0.424 0.914 0.330 0.889 0.548 0.641
 Llama-3.3-70B-Instruct 0.862 0.391 0.905 0.585 0.915 0.400 0.916 0.718 0.712
-QwQ-32B 0.892 0.663 0.922 0.716 0.922 0.570 0.933 0.793 0.801
 GLM-4-32B-0414 0.855 0.347 0.897 0.512 0.909 0.325 0.909 0.664 0.677
@@
 Baselines.
-We choose state-of-the-art general and evaluation-specific LLMs as our baselines.
-(1) General LLMs: We adopt 4 advanced proprietary LLMs, GLM-4-Plus (GLM et al.
-2024), GPT-4.1 (OpenAI 2023), Deepseek-R1 (Guo et al.
-2025a) and o4-mini (Jaech et al.
-2024), as well as 4 strong open-source LLMs, GLM-4-32B-0414 (GLM et al.
-2024), QwQ-32B (Team 2025), LLama-3.3-70B-Instruct (Grattafiori et al.
-2024), and Qwen2.5-72B-Instruct (Yang et al.
-2024) as our general baselines.
-These models can serve as evaluators for instruction-following via elaborate prompts without further training.
-We explore two prompting strategies: 1) Checklist-Level Prompt: requires the LLMs to evaluate the following of all constraints in the given checklist within a single inference pass, which is the same as IF-Critic.
-2) Constraint-Level Prompt: requires the LLMs to evaluate each constraint individually.
-(2) Evaluation-specific LLMs: We select Skywork-Reward-V2-Llama-3.1-8B-40M (Liu et al.
-2025a), Prometheus-BGB-8x7B (Kim et al.
-2024b), RM-R1-Deepseek-Distilled-Qwen-32B (Chen et al.
-2025), and RRM-32B (Guo et al.
-2025b) as our evaluation-specific baselines.
-These models are mainly designed for pairwise comparison between two responses.
+We choose state-of-the-art general and evaluation-specific LLMs as our baselines, which are listed in Appendix C.
+General LLMs can serve as evaluators for instruction-following via elaborately prompting.
+We explore two prompting strategies: (1) Checklist-Level Prompt: evaluate all constraints in the given checklist within a single inference pass, which is the same as IF-Critic.
+(2) Constraint-Level Prompt: evaluate each constraint individually.
+Evaluation-specific LLMs are mainly designed for pairwise comparison.
 To accommodate this evaluation setting, we convert the above datasets into pairwise versions.
-Specifically, for two responses generated from the same instruction, we select pairs in which one response follows all the constraints while the other does not to construct pairwise samples.
-The rewards determined by the critiques of IF-Critic according to Equation 6 for both responses are converted into pairwise comparisons to calculate the agreement rate.
-To ensure a fair comparison, we remove any samples that result in a tie for IF-Critic.
+Specifically, for two responses generated from the same instruction, we select pairs in which one response follows all constraints while the other does not to construct pairwise samples.
+The rewards of IF-Critic according to Equation 6 for both responses are converted into pairwise comparisons to calculate the agreement rate.
+To ensure a fair comparison, we remove samples that result in a tie for IF-Critic.
 Greedy search decoding is used for reproducibility.
 Implementation Details.
-We choose Qwen2.5-14B-Instruct (Yang et al.
-2024) as the base model for both the checklist generator and IF-Critic.
-Deepseek-R1 is required to generate $N=5$ expert critiques per response.
-We use difflab [^3] https://docs.python.org/3/library/difflib.html to implement the text similarity measure $u(\cdot)$ for final explanation selection.
+We choose Qwen-2.5-14B-Instruct Yang et al.
+(2024) as the base model for both the checklist generator and IF-Critic.
+We also experiment with other base models, and the results are provided in Appendix E.
+Deepseek-R1 generates $N=5$ expert critiques per response.
 During IF-Critic training, the ratio of $D_{\text{sft}}$ to $D_{\text{ref}}$ is set to 6:4.
 The number of self-sampled critiques, $M$, is set to 10, and up to 1 preference pair is constructed for each evaluation input.
-More details of IF-Critic training are in the Appendix.
+More implementation details are in Appendix D.
 Model EvalCritic CFBench TRACE Multi-IF
@@
 w/o Preference Learning 0.815 0.810 0.810 0.841
-Table 3: The average of Positive F1 and Negative F1 under different ablation settings for IF-Critic.
+Table 3: The average of Positive F1 and Negative F1 scores under different ablation settings for IF-Critic.
 Policy Model Method Reward Model / Critic Relative Multi-IF CFBench SysBench
@@
 Table 4: Experimental results (%) on Multi-IF, CFBench, and SysBench.
+The best performance among different critics is bold.
 Relative Time refers to the relative time to IF-Critic of reward calculation in DPO and average per-step training in GRPO.
-Skywork-Reward-V2-Llama-3.1-8B-40M is a reward model that generates scalar rewards, whereas QwQ-32B and IF-Critic generate natural language critiques.
+Skywork-Reward-V2 is a reward model that generates scalar rewards, whereas QwQ-32B and IF-Critic generate natural language critiques.
 Main Results.
 Table 1 presents the F1-scores for both positive and negative classes in the instruction-following evaluation task of general LLMs, while Table 2 shows the pairwise agreement for evaluation-specific LLMs.
-Firstly, IF-Critic demonstrates superior performance on various evaluation benchmarks, outperforming all baselines on average F1.
-Notably, it surpasses the expert critiques generation models, Deepseek-R1, on all benchmarks, and o4-mini on three out of four benchmarks.
+Firstly, IF-Critic demonstrates strong performance on various benchmarks, outperforming all baselines on average F1, including o4-mini and Gemini-3-Pro.
 In contrast, the performance of existing general LLMs remains limited, with relatively low negative F1.
@@
 Secondly, existing evaluation-specific LLMs still perform poorly in evaluating instruction-following and lag far behind IF-Critic, highlighting the necessity of a critic specifically designed for this task.
-Finally, reasoning models (i.e., o4-mini, Deepseek-R1, and QwQ-32B) perform better with Checklist-Level Prompt, while other LLMs benefit more from Constraint-Level Prompt.
-We attribute this to two factors: (1) Evaluating multiple constraints at one inference pass increases task complexity.
-(2) With the assistance of the constraint checklist, long-chain reasoning enables a more holistic perception of the relationships between constraints.
-This validates the rationale of our checklist-guided critique generation paradigm.
+Finally, reasoning models (e.g, o4-mini, QwQ-32B) typically perform better with Checklist-Level Prompt.
+This may be attributed that long-chain reasoning enables a more holistic perception of the relationships between constraints with the help of checklists.
+This observation validates the rationale of our checklist-guided critique generation paradigm.
 Ablation Study.
-We conduct ablation studies on three main factors of our framework: training data format, critique filtering mechanism, and preference optimization method.
-For the training data format, we decompose the expert critiques to train an alternative model that generates critiques for individual constraints.
-Table 3 shows that the performance of IF-Critic declines across all benchmarks, demonstrating that checklist-guided training can facilitate the perception of instructions and contribute to the final performance.
-For the critique filtering mechanism, we remove each step individually to explore its impact on evaluation performance.
-Table 3 shows that all steps contribute to the final performance, with the removal of rule-augmented verification yielding the largest performance drop, underscoring its key role in improving the discriminability for length constraints.
-Notably, removing the entire mechanism and directly using all raw training data leads to the most severe performance decline, highlighting the effectiveness of critique filtering.
-For the preference optimization method, we explored three variants of our method: (1) Constructing preference pairs directly based on the agreement rates between the expert critiques, denoted as Vanilla DPO; (2) Replacing misaligned segments of rejected critiques with expert critiques rather than the best self-sampled critiques; and (3) Performing SFT on the entire critique training dataset without preference optimization.
-Table 3 demonstrates that our constraint-level preference optimization method achieves the best performance, evidencing its effectiveness in enabling better alignment with expert critiques.
+To investigate the impact of each factor of our framework, we conduct additional ablation studies.
+For data format, we decompose expert critiques to train an alternative model that generates critiques for individual constraints.
+Table 3 shows that the performance of IF-Critic declines across all benchmarks, demonstrating that checklist-guided training can facilitate instruction perception and contribute to the final performance.
+For critique filtering mechanism, we remove each step individually to explore its impact on evaluation performance.
+Table 3 shows that all steps contribute to the final performance, but removing rule-augmented verification yields the largest performance drop, underscoring its key role in improving discriminability for length constraints.
+Notably, directly using all raw training data without critique filtering leads to the most severe performance decline, highlighting its effectiveness.
+For preference optimization method, we explored three variants of our method: (1) Constructing preference pairs directly based on the agreement rates between expert critiques, denoted as Vanilla DPO, (2) Replacing misaligned segments of rejected critiques with expert critiques rather than the best self-sampled ones, and (3) Performing SFT on the entire training dataset without preference optimization.
+Table 3 shows that our constraint-level preference optimization method achieves the best performance, evidencing its effectiveness in enabling better alignment with expert critiques.
 Figure 3: Explanation quality evaluation results. The percentages indicate the preference between IF-Critic and other evaluation models via human annotation.
@@
 As shown in Figure 3, IF-Critic surpasses QwQ-32B and Deepseek-R1 in win rates by 9.3% and 7.7%, respectively, and performs on par with o4-mini.
+Many results are ties because all four evaluators are strong and correct on many constraints.
 This demonstrates that IF-Critic can generate high-quality explanations for instruction-following evaluation.
-Instruction-Following Optimization
+4.2 Instruction-Following Optimization
 Training Datasets & Settings.
-To collect diverse and high-quality complex instructions for model training, we first curate a set of seed instructions from real application scenarios.
+To collect diverse and high-quality complex instructions for model training, we first curate seed instructions from real-world application scenarios.
 These scenarios are distinct from the training data of IF-Critic and generally exhibit lower instruction complexity.
-Stratified sampling based on task categories and LLM-based quality filtering is adopted to enhance their diversity and quality.
-Subsequently, we employ Deepseek-R1 to introduce constraints to these seed instructions, guided by a comprehensive constraint taxonomy encompassing 6 primary constraint categories and 4 constraint composition types, thereby increasing their complexity.
-Deepseek-R1 is also leveraged to validate the generated instructions, ensuring they are reasonable, unambiguous, and logically consistent.
-Further details are in the Appendix.
+Stratified sampling based on task categories is adopted to enhance diversity.
+Subsequently, we employ Deepseek-R1 to introduce constraints to these instructions based on a comprehensive constraint taxonomy that encompasses 6 primary constraint categories and 4 constraint composition types.
+Deepseek-R1 is also used to validate the generated instructions, ensuring they are reasonable, unambiguous, and logically consistent.
 In total, we obtain 18.5k high-quality complex instructions for training.
+Further details are in Appendix F.
 We choose Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct as the initial policy models for DPO and GRPO training.
-Three evaluation models are used to provide reward signals: IF-Critic, the state-of-the-art reward model Skywork-Reward-V2-Llama-3.1-8B-40M, and open-source general LLM QwQ-32B.
+Three evaluation models are used to provide reward signals: IF-Critic, the state-of-the-art reward model Skywork-Reward-V2-Llama-3.1-8B-40M, and the powerful open-source general LLM QwQ-32B.
 For DPO training, we sample 10 responses per instruction, with a temperature of 1.0 and a top-p value of 0.9.
 The responses with the highest and lowest rewards are used to construct preference pairs.
-For GRPO training, we perform 32 rollouts per instruction for value estimation.
-More training details can be found in the Appendix.
+For GRPO training, we perform 32 rollouts per instruction.
+More training details are in Appendix F.
 Evaluation Benchmarks.
-We evaluate our models on three popular instruction-following benchmarks: Multi-IF (He et al.
-2024b), which assesses multi-turn and multilingual hard constraints following; CFBench (Zhang et al.
-2024a), which covers a comprehensive range of constraint types; and Sysbench (Qin et al.
-2025), which focuses on the following of system prompts.
+We evaluate our models on three popular instruction-following benchmarks: Multi-IF He et al.
+(2024b), which assesses multi-turn and multilingual hard constraints following; CFBench Zhang et al.
+(2024a), which covers a comprehensive range of constraint types; and Sysbench Qin et al.
+(2025), which focuses on the following of system prompts.
 GPT-4o-2024-11-20 is used as the evaluation model for CFBench and SysBench.
-For Multi-IF, we follow RPO (Huang et al.
-2025) and report the average performance across the English and Chinese subsets.
+For Multi-IF, we follow RPO Huang et al.
+(2025) and report the average performance across the English and Chinese subsets.
 Main Results.
 The main results are presented in Table 4.
-Firstly, optimization with IF-Critic brings substantially superior performance gains across various benchmarks compared to all strong baselines, substantiating its advanced capacity of reliable instruction-following evaluation.
+Firstly, optimization with IF-Critic brings substantially superior performance gains across various benchmarks compared to all strong baselines, substantiating its advanced capacity of instruction-following evaluation.
 Notably, LLMs after GRPO training with IF-Critic even achieve comparable performance to significantly larger LLMs (32B or 70B) of the same model family.
-In contrast, the general reward model, Skywork-Reward-V2-Llama-3.1-8B-40M, is not well-suited for instruction-following optimization, yielding only marginal improvements.
-Secondly, compared to the powerful baseline LLM critic, QwQ-32B, IF-Critic only incurs less than a third of computational overhead, highlighting its practicality.
-Finally, between the two commonly used optimization methods, DPO and GRPO, GRPO proves more effective for instruction-following tasks, delivering greater performance gains.
-The reward curves during GRPO training are provided in the Appendix.
-We also validate that optimization with IF-Critic does not compromise the general performance of LLMs in the Appendix.
-Conclusion
-In this work, we propose IF-Critic, a LLM critic which is capable of providing fine-grained assessments for instruction-following.
-IF-Critic adopts a checklist-guided critique generation paradigm that evaluates the following of all constraints in a given checklist within one inference pass, improving its efficiency and reliability.
-To construct IF-Critic, we collect high-quality critique training data through a multiple-stage critique filtering mechanism and employ a constraint-level preference optimization method to train IF-Critic.
-Extensive experiments demonstrate that IF-Critic can beat strong LLM-as-a-Judge baselines and bring substantially superior performance gains to LLMs in instruction-following optimization, compared to strong LLM critic baselines.
-Appendix A Details of IF-Critic Training
-We use the Llama-Factory (Zheng et al.
-2024) framework and the Zero Redundancy Optimizer (ZeRO) (Rajbhandari et al.
-2020) stage 3 from the Deepspeed library (Rasley et al.
-2020) to train IF-Critic.
+In contrast, the general reward model, Skywork-Reward-V2, is ill-suited for instruction-following optimization, yielding only marginal improvements.
+Secondly, compared to the powerful LLM critic, QwQ-32B, IF-Critic only incurs less than a third of computational overhead, highlighting its practicality.
+Finally, GRPO proves more effective for instruction-following tasks compared to DPO, delivering greater performance gains.
+The reward curves during GRPO training are provided in the Appendix G.
+We also validate that optimization with IF-Critic does not compromise the general performance of LLMs in the Appendix H.
+5 Conclusion
+Our work proposes IF-Critic, an LLM critic for fine-grained instruction-following evaluation.
+IF-Critic adopts a checklist-guided critique generation paradigm that evaluates all constraints in a checklist within one inference pass, improving efficiency and reliability.
+We collect high-quality critique training data through a multi-stage filtering mechanism and employ constraint-level preference optimization to train IF-Critic.
+Extensive experiments show that IF-Critic has stronger evaluation ability and can bring superior performance gains to LLMs in instruction-following optimization, compared to powerful LLM critic baselines.
+6 Limitations
+The limitations of our work are summarized as follows:
+Scope of Rule-Augmented Verification.
+In the critique training data construction stage of IF-Critic, we only introduce rule-augmented verification for length-related constraints, while other code-verifiable constraints (e.g., keywords or structural format) are not comprehensively covered.
+We prioritize length-related constraints because they are prevalent in real-world complex instructions, and the previous version of IF-Critic performs well on other verifiable constraints but struggles with length constraints, as observed in our primary experiments.
+Although this approach has yielded substantial performance gains, extending this approach to a broader spectrum of constraint types remains an important direction for future work.
+Potential Evaluation Bias.
+Similar to other model-based evaluation methods, IF-Critic may also suffer from potential evaluation biases such as self-enhancement or verbosity bias Zheng et al.
+(2023).
+These biases could harm the evaluation correctness of certain models.
+Although our multi-stage critique filtering mechanism has effectively mitigated noise and bias in training critique collection, incorporating inference-time strategies (e.g., multi-agent debate) may further attenuate bias and enhance evaluation performance.
+We reserve further investigation of evaluation bias mitigation as important future work.
+Ethical Considerations
+In this work, we recruit some human annotators to validate the quality of training data and the performance of IF-Critic.
+The annotator pool primarily consists of college students.
+Throughout the data annotation process, we adhere to several key principles.
+Firstly, all annotators are fully informed about the purpose of our study and the involved tasks.
+Secondly, all annotators received fair compensation for their contributions based on the market price.
+Finally, any harmful instructions are filtered before annotation to avoid ethical issues.
+Appendix A Model List for Response Generation
+We use 15 representative LLMs for response generation, including GPT-4o OpenAI (2023), GPT-4o-mini OpenAI (2023), seven versions of GLM-4 GLM et al.
+(2024), three versions of Qwen2.5 Yang et al.
+(2024), Deepseek-R1 Guo et al.
+(2025a), Doubao1.5 Doubao Team (2025), and Claude3.7-sonnet Anthropic (2023).
+Appendix B List for Prompt Templates
+This section lists all the prompt templates applied throughout this work, including the prompt for scoring the quality of user instruction in Table 8, the prompt for constraint checklist generation in Table 9, the prompt for checklist-guided critique generation in Table 10, the prompt for cross-model verification in Table 11, and the prompt for rule-augmented verification in Table 12 and 13.
+Appendix C List for Baseline Evaluation Models
+For general LLMs, we adopt 3 advanced proprietary LLMs, GPT-4.1 OpenAI (2023), o4-mini Jaech et al.
+(2024), and Gemini-3-Pro Team et al.
+(2023), as well as 5 strong open-source LLMs, GLM-4-32B-0414 GLM et al.
+(2024), LLama-3.3-70B-Instruct Grattafiori et al.
+(2024), Qwen2.5-72B-Instruct Yang et al.
+(2024), QwQ-32B Team (2025), and Deepseek-R1 Guo et al.
+(2025a) as our general baselines.
+For evaluation-specific LLMs, we select 4 powerful discriminative or generative reward models, Skywork-Reward-V2-Llama-3.1-8B-40M Liu et al.
+(2025a), Prometheus-BGB-8x7B Kim et al.
+(2024b), RM-R1-Deepseek-Distilled-Qwen-32B Chen et al.
+(2025), and RRM-32B Guo et al.
+(2025b) as our evaluation-specific baselines.
+Appendix D Implementation Details of IF-Critic
+We use the Llama-Factory Zheng et al.
+(2024) framework and the Zero Redundancy Optimizer (ZeRO) Rajbhandari et al.
+(2020) stage 3 from the Deepspeed library Rasley et al.
+(2020) to train IF-Critic.
 All experiments are conducted on 8 H800 GPUs.
-We utilize the AdamW (Kingma and Ba 2014) optimizer with a weight decay of 0.1, and set the maximum sequence length to 8192 tokens during all the experiments.
-The overall API cost for critique training data collection is approximately 4000\$.
+We utilize the AdamW Kingma and Ba (2014) optimizer with a weight decay of 0.1, and set the maximum sequence length to 8192 tokens during all the experiments.
 During the supervised fine-tuning, the peak learning rate is set to 5e-6 with a 10% warmup ratio and a linear scheduler.
@@
 Training utilizes a sigmoid loss function with a beta value of 0.1 and spans 1 epoch.
-Following previous works (Hou et al.
-2024), an additional SFT loss is added to the chosen critique with a weight of 0.1.
-Appendix B Details of Training Instructions Collection in Instruction-Following Optimization
+Following previous works Hou et al.
+(2024), an additional SFT loss is added to the chosen critique with a weight of 0.1.
+The overall API cost for critique training data collection is approximately 4000\$.
+During final explanation selection, we use difflab [^2] https://docs.python.org/3/library/difflib.html to implement the text similarity measure $u(\cdot)$.
+Appendix E Experimental Results on Other Base Models
+Base Model EvalCritic CFBench TRACE Multi-IF
+Qwen2.5-14B-Insturct 0.861 0.863 0.840 0.895
+Qwen-2.5-7B-Instruct 0.822 0.798 0.806 0.813
+Qwen-3-8B 0.855 0.823 0.820 0.848
+GLM-4-9B-0414 0.797 0.810 0.801 0.810
+Qwen-3-14B 0.868 0.863 0.828 0.860
+Qwen-2.5-32B-Instruct 0.879 0.854 0.833 0.867
+Table 5: The average of Positive F1 and Negative F1 scores of IF-Critic under different base models. Qwen-2.5-14B-Instruct is the base model we used for the main experiments.
+In our primary experiment, we experiment with various base models for IF-Critic, and the results are presented in Table 5.
+We observe that LLMs with approximately 8B parameters are still insufficient to fully acquire critique generation capabilities, exhibiting a notable performance gap to larger LLMs.
+In contrast, LLMs with approximately 14B parameters are generally adequate for learning critique generation, and further increasing model scale yields negligible performance improvement.
+Considering the trade-off between performance and efficiency, we ultimately choose Qwen-2.5-14B-Instruct as the base model for our main experiments.
+Appendix F Details of Instruction-Following Optimization
+Training Instructions Collection.
 To ensure the diversity and complexity of constraints, we first construct a comprehensive and hierarchical constraint taxonomy, which consists of 6 primary categories (i.e., format constraints, content constraints, behavioral constraints, style constraints, role constraints, and environmental constraints) and 54 secondary categories.
 During constraint synthesis, we randomly sample 2 to 5 secondary constraint categories and require Deepseek-R1 to introduce several corresponding constraints into the seed instructions, thereby increasing their complexity.
-In contrast to previous work that only considers independent and atomic constraints (Jiang et al.
-2023; Ren et al.
-2025; Peng et al.
-2025), we further ask Deepseek-R1 to compose multiple constraints according to the composition types proposed in ComlexBench (Wen et al.
-2024), including: (1) And: Coordination between different constraints, (2) Chain: Sequential completion of constraints, (3) Selection: Conditional selection of constraints, and (4) The nested structure of the above composition types.
+In contrast to previous work that only considers independent and atomic constraints Jiang et al.
+(2023); Ren et al.
+(2025); Peng et al.
+(2025), we further ask Deepseek-R1 to compose multiple constraints according to the composition types proposed in ComlexBench Wen et al.
+(2024), including: (1) And: Coordination between different constraints, (2) Chain: Sequential completion of constraints, (3) Selection: Conditional selection of constraints, and (4) The nested structure of the above composition types.
 In this way, our constraint synthesis can more comprehensively encompass the complex instruction types in real-world scenarios.
-In this scenario, it is notable that most of the hard constraints cannot be reliably verified using automatic verification codes alone, such as constraints targeting for specific segments of the response (e.g., Generate 10 titles, each no more than 8 words) and the composition of soft and hard constraints (e.g., Bold all adjectives that express feelings, such as **happy**), as it is challenging for verification code to accurately extract the corresponding segments of the response that should follow these constraints.
+In this scenario, it is notable that most of the hard constraints cannot be reliably verified using automatic verification codes alone, such as constraints targeting for specific segments of the response (e.g., Generate 10 titles, each no more than 8 words) and the composition of soft and hard constraints (e.g., Bold all adjectives that express feelings, such as happy), as it is challenging for verification code to accurately extract the corresponding segments of the response that should follow these constraints.
 We provide some examples of the seed instructions and the corresponding final constructed complex instructions in Table 6.
@@
 Seed Instruction Write a conversation between a Chinese international student and a Japanese student who decide to try a new restaurant together.
-While they’re having their meal, one of them suggests that they go to see a movie afterward.
+While they’re having their meal, one of them suggests that they go to see a movie afterwards.
 Final Instruction Please write a conversation between Xiaolin, a Chinese international student, and Yamada, a Japanese student.
 In the conversation, they decide to try a new restaurant together.
-While they’re having their meal, one of them suggests that they go to see a movie afterward.
+While they’re having their meal, one of them suggests that they go to see a movie afterwards.
 Ensure the following requirements are satisfied:
@@
 3.
-Xiaolin should naturally mention Mid-Autumn Festival customs in the conversation, while Yamada should incorporate elements of the Obon Festival.
-Please use specific festival details (e.g., types of mooncakes or forms of Bon Odori).
+Xiaolin should naturally mention Mid-Autumn Festival customs in the conversation while Yamada should incorporate elements of the Obon Festival.
+Please use specific festival details (e.g., types of mooncakes, or forms of Bon Odori).
 4. Please write the conversation in Chinese. Indicate the speaker (Xiaolin: / Yamada:) at the start of each turn. The entire conversation should be no more than 10 turns.
@@
 Table 6: Some examples of the seed instructions and the corresponding final constructed complex instruction in the instruction-following optimization.
-Appendix C Details of Instruction-Following Optimization
-For DPO training, we use the Llama-Factory (Zheng et al.
-2024) framework, and all other settings are the same as the preference optimization for IF-Critic.
-During the reward calculation, IF-Critic and QwQ-32B utilize the vllm (Kwon et al.
-2023) inference framework, while Skywork-Reward-V2-Llama-3.1-8B-40M is deployed by sglang [^4] https://github.com/sgl-project/sglang (since this reward model is not compatible with the vllm framework currently).
+Training Configuration.
+For DPO training, we use the Llama-Factory Zheng et al.
+(2024) framework, and all other settings are the same as the preference optimization for IF-Critic.
+During the reward calculation, IF-Critic and QwQ-32B utilize the vllm Kwon et al.
+(2023) inference framework, while Skywork-Reward-V2-Llama-3.1-8B-40M is deployed by sglang [^3] https://github.com/sgl-project/sglang (since this reward model is not compatible with the vllm framework currently).
 The reward calculation and DPO training are both conducted on 8 H800 GPUs.
-For GRPO training, we use the VeRL (Sheng et al.
-2025) framework.
+For GRPO training, we use the VeRL Sheng et al.
+(2025) framework.
 The learning rate is set to 2e-6, with the maximum sequence length for both prompts and responses set to 4096 tokens.
 The train-batch size and PPO mini-batch size are each set to 32.
-For rollout responses generation, we employ the vllm (Kwon et al.
-2023) framework, utilizing 60% of the available GPU memory, with a temperature and top-p value of 1.0.
+For rollout responses generation, we employ the vllm Kwon et al.
+(2023) framework, utilizing 60% of the available GPU memory, with a temperature and top-p value of 1.0.
 To encourage stable training, we also incorporate KL divergence regularization with a coefficient of 0.001, using the low-variance KL implementation.
 We save checkpoints every 30 steps during training.
-Following TULU 3 (Lambert et al.
-2024) and VerIF (Peng et al.
-2025), we use IFEval (Zhou et al.
-2023) as the validation set to select the best checkpoint.
+Following TULU 3 Lambert et al.
+(2024) and VerIF Peng et al.
+(2025), we use IFEval Zhou et al.
+(2023) as the validation set to select the best checkpoint.
 All experiments are conducted on 24 H800 GPUs, with 8 GPUs dedicated to model training and the remaining 16 GPUs allocated for API deployment of the instruction-following critic or reward model.
@@
 Skywork-Reward-V2 refers to Skywork-Reward-V2-Llama-3.1-8B-40M.
-Appendix D Reward Curves During GRPO Training
+Appendix G Reward Curves During GRPO Training
 We provide the reward curves during GRPO training when IF-Critic and QwQ-32B are employed as the LLM critic in Figure 4, and Skywork-Reward-V2-Llama-3.1-8B-40M as the reward model in Figure 5.
@@
 Using it as a reward model for GRPO training significantly increases the average response length of the policy model, resulting in a longer average per-step training time compared to IF-Critic.
-Appendix E Experimental Results on General Benchmarks
-We provide the performance of LLMs after instruction-following optimization in four general benchmarks: AlignBench (Liu et al.
-2024), MMLU-Pro (Wang et al.
-2024b), Omni-MATH (Gao et al.
-2025), and HumanEval (Chen et al.
-2021) in Table 6.
+Appendix H Experimental Results on General Benchmarks
+We provide the performance of LLMs after instruction-following optimization in four general benchmarks: AlignBench Liu et al.
+(2024), MMLU-Pro Wang et al.
+(2024), Omni-MATH Gao et al.
+(2025), and HumanEval Chen et al.
+(2021) in Table 7.
 GPT-4o-2024-11-20 is used as the evaluation model for AlignBench, and we use OmniJudge [^4] https://huggingface.co/KbsdJames/Omni-Judge for the evaluation of Omni-MATH.
 The results reveal that conducting instruction-following optimization with IF-Critic does not harm the general performance of LLMs.
-Appendix F List of Prompt Templates
-This section lists all the prompt templates applied throughout this work, including the prompt for scoring the quality of user instruction in Table 7, the prompt for constraint checklist generation in Table 8, the prompt for checklist-guided critique generation in Table 9, the prompt for cross-model verification in Table 10, and the prompt for rule-augmented verification in Table 11 and 12.
-Appendix G Human Annotation Guideline for Constraint Verification
+Appendix I Human Annotation Guideline for Constraint Verification
 The human annotation guideline for constraint verification in the section Instruction-Following Evaluation is shown in Table 14.
@@
 ## Evaluation Criteria:
-1. Low Quality: The user instruction contains significant issues, such as unclear, incomplete, or ambiguous information, making it impossible to determine the intent behind the instruction.
+1. Low Quality: The user instruction contains significant issues such as unclear, incomplete, or ambiguous information, making it impossible to determine the intent behind the instruction.
 2. Medium Quality: The user instruction may include some unclear, incomplete, or ambiguous elements; however, the overall intent can still be inferred.
@@
 (1) Each constraint you output must be a verbatim excerpt from the given instruction. You must not fabricate, paraphrase, or add any content.
-(2) You must extract and output all constraints in the order in which they appear in the instruction, without omitting any or inventing constraints not present in the instruction.
+(2) You must extract and output all constraints in the order in which they appear in the instruction, without omitting any constraints or inventing constraints not present in the instruction.
 To reiterate: Constraints refer to all instructional content, excluding auxiliary information such as background knowledge, text materials, and in-context examples.
@@
 Please think carefully and provide a detailed analysis of whether the AI assistant’s response follows each item on the constraint checklist.
-You need to analyze and assess every constraint in the checklist, and you do not need to analyze anything that is not listed in the checklist.
+You need to analyze and assess every constraint in the checklist, and do not need to analyze anything that is not listed in the checklist.
 You must output your explanation and judgment for each constraint, strictly following the format below:
@@
 (2) Your judgment should be as strict as possible.
-You should output ”[[The AI assistant’s response follows this constraint]]” only if the response completely follows every part of the constraint as stated, without any omission or mistake.
+You should output "[[The AI assistant’s response follows this constraint]]" only if the response completely follows every part of the constraint as stated, without any omission or mistake.
 (3) When analyzing a given constraint, you should not take into account whether other constraints are followed.
@@
 We will provide you with a user instruction, an AI assistant’s response, a specific constraint from the user instruction, and a critique of the AI’s response regarding whether the response follows this constraint from the user instruction.
-The critique comprises two parts: ”Explanation” and ”Judgment”.
+The critique comprises two parts: "Explanation" and "Judgment".
 Your task is to carefully analyze this critique to judge whether it is reasonable and correct.
-After your analysis, output either ”[[The given critique is correct]]” or ”[[The given critique is not correct]]” as your final conclusion.
+After your analysis, output either "[[The given critique is correct]]" or "[[The given critique is not correct]]" as your final conclusion.
 A correct critique must simultaneously satisfy all of the following three principles:
@@
 (2) Be as strict as possible.
-Output ”[[The AI assistant’s response follows this requirement]]” only if the response completely follows every part of the requirement as stated, without any omission or mistake.
+Output "[[The AI assistant’s response follows this requirement]]" only if the response completely follows every part of the requirement as stated, without any omission or mistake.
 (3) Independently analyzes the given constraint, and does not take into account whether other constraints are followed.
@@
 You only need to judge whether the explanation and judgment within the critique are logically consistent, specifically:
-(1) If the explanation section determines that the response follows the constraint, the judgment should be: ”[[The AI assistant’s response follows the constraint]]”.
-(2) If the explanation section identifies any mistake or shortcoming in the response, indicating that the constraint is not fully followed, the judgment should be: ”[[The AI assistant’s response does not follow the constraint]]”.
+(1) If the explanation section determines that the response follows the constraint, the judgment should be: "[[The AI assistant’s response follows the constraint]]".
+(2) If the explanation section identifies any mistake or shortcoming in the response, indicating that the constraint is not fully followed, the judgment should be: "[[The AI assistant’s response does not follow the constraint]]".
 Remember the following points:
-(1) There should not be situations such as ”partially follow the constraint”.
-As long as the explanation identifies any mistakes or shortcomings in the response that result in the constraint not being fully followed, the judgment should be: ”[[The AI assistant’s response does not follow the constraint]].”
-(2) If the explanation states that the constraints and the response are unrelated and there is no need for evaluation, the judgment should be: ”[[The AI assistant’s response follows the constraint]].”
+(1) There should not be situations such as "partially follow the constraint".
+As long as the explanation identifies any mistakes or shortcomings in the response that result in the constraint not being fully followed, the judgment should be: "[[The AI assistant’s response does not follow the constraint]]."
+(2) If the explanation states that the constraints and the response are unrelated and there is no need for evaluation, the judgment should be: "[[The AI assistant’s response follows the constraint]]."
 Please analyze the logical consistency between the explanation and judgment following the above principles.
-Output your analysis first, and then output either ”[[Explanation and Judgment are consistent]]” or ”[[Explanation and Judgment are not consistent]]” to represent your final verdict.
+Output your analysis first, and then output either "[[Explanation and Judgment are consistent]]" or "[[Explanation and Judgment are not consistent]]" to represent your final verdict.
 Here are the constraint and the critique to be judged:
@@
 We will provide you with a user instruction, an AI assistant’s response, and a specific constraint from the user instruction.
-Please first determine whether the given constraint is related to length (e.g., ”The article should be no less than 800 words” or ”Each title must be 8 characters long”).
+Please first determine whether the given constraint is related to length (e.g., "The article should be no less than 800 words" or "Each title must be 8 characters long").
 If so, please extract the segments from the AI assistant’s response that should follow this length requirement (extract the original text from the response without any modifications).
@@
 1. (If the constraint is NOT related to length)
-”””json
-   ”Length Constraint” : False
+"""json
+   "Length Constraint" : False
 2. (If the constraint is related to length)
-”””json
-   ”Length Constraint” : True,
-   ”Extracted Segments” : [
-       ”Length Requirement within the Constraint”: …(Provide the length requirement within the constraint. Be sure to extract the original text from the constraint without making any modification),
-       ”Corresponding Segment in Response”: …(Provide the segment of the response that should follow this requirement without making any modification.
-If no corresponding segment exists in the response, output ”No corresponding segment exists.”)
-       ”Length Requirement within the Constraint”: …,
-       ”Corresponding Segment in Response”: …
+"""json
+   "Length Constraint" : True,
+   "Extracted Segments" : [
+       "Length Requirement within the Constraint": …(Provide the length requirement within the constraint. Be sure to extract the original text from the constraint without making any modification),
+       "Corresponding Segment in Response": …(Provide the segment of the response that should follow this requirement without making any modification.
+If no corresponding segment exists in the response, output "No corresponding segment exists.")
+       "Length Requirement within the Constraint": …,
+       "Corresponding Segment in Response": …
      … (Analogously, extracting all segments in the response that should follow the length requirements of the given constraint)
@@
 Each item consists of three elements: (1) a length requirement within the constraint, (2) a segment of the responses that should follow the length requirement, and (3) the actual length of the segment.
-It is important that you consider the ”actual length provided in the JSON data” as the golden-truth length, which is absolutely accurate.
+It is important that you consider the "actual length provided in the JSON data" as the golden-truth length, which is absolutely accurate.
 Do not make your own calculation about “the length of the segment”.
-You must rely solely on the ”actual length” as the basis for length validation.
+You must rely solely on the "actual length" as the basis for length validation.
 Output your explanation and judgment for the given constraint based on the actual length provided in the JSON data, strictly following the format below:
@@
 (2) Your judgment should be as strict as possible.
-You should output ”[[The AI assistant’s response follows this constraint]]” only if the response completely follows every part of the constraint as stated, without any omission or mistake.
-(3) Independently analyzing the given constraint, without taking into account whether other constraints are followed.
+You should output "[[The AI assistant’s response follows this constraint]]" only if the response completely follows every part of the constraint as stated, without any omission or mistake.
+(3) Independently analyzing the given constraint, and does not take into account whether other constraints are followed.
 (4) The given constraint may contain both length requirements and other requirements.
-For length requirements, you must rely solely on the ”actual length provided in the JSON data” as the basis for length validation.
+For length requirements, you must rely solely on the "actual length provided in the JSON data" as the basis for length validation.
 For other requirements, please refer to the above principles and carefully evaluate them by yourself.
-(5) When generating your explanation, it is forbidden to mention phrases such as ”according to the given JSON data” or ”based on the provided actual length”.
+(5) When generating your explanation, it is forbidden to mention phrases such as "according to the given JSON data" or "based on the provided actual length".
 Your explanation should convey that the length is calculated independently by you, without referencing the given JSON data.
@@
 The checklist contains some of the constraints within the instruction.
-Your task is to verify whether the model response follows each constraint in the checklist, choose ”Followed” or ”Not Followed”, and provide a brief justification.
+Your task is to verify whether the model response follows each constraint in the checklist, choose "Followed" or "Not Followed", and provide a brief justification.
 Task Details:
@@
 Your judgments must be strict.
-Only responses that fully satisfy a constraint can be judged as ”Followed”.
-If there is any omission or error regarding a constraint, it must be judged as ”Not Followed”.
+Only responses that fully satisfy a constraint can be judged as "Followed".
+If there is any omission or error regarding a constraint, it must be judged as "Not Followed".
 3. Please focus exclusively on the constraints within the given checklist. It is unnecessary to consider whether the response follows any other constraints beyond the checklist.